### v3

# Representation theorems for dynamic convex risk measures

Abstract
In this paper, we prove that under the domination condition:
${\cal{E}}^{-\mu,-\nu}[-\xi|{\cal{F}}_{t}]\leqslant\rho_{t}(\xi)\leqslant{\cal{E}}^{\mu,\nu}[-\xi|{\cal{F}}_{t}],\quad\forall\xi\in\mathcal{L}^{\exp}_{T}\ (\text{resp.}\ L^{2}(\mathcal{F}_{T})),\ \forall t\in[0,T],$
where ${\cal{E}}^{\mu,\nu}$ is the $g$-expectation with generator $\mu|z|+\nu|z|^{2},\mu\geqslant 0,\nu\geqslant 0$, the dynamic convex (resp. coherent) risk measure $\rho$ admits a representation as a $g$-expectation, whose generator $g$ is convex (resp. sublinear) in the variable $z$ and has a quadratic (resp. linear) growth. As an application, we show that such dynamic convex (resp. coherent) risk measure $\rho$ admits a dual representation, where the penalty term (resp. the set of probability measures) is characterized by the corresponding generator $g$.
Keywords: Convex risk measures; Coherent risk measures; Backward stochastic differential equation; $g$-expectation; Dual representation
AMS Subject Classification: 60H10
1 Introduction
An important task in risk management is to quantify the risks of financial positions. Artzner et al. [1] introduced the notion of coherent risk measures, which are defined by a set of four desirable properties for measures of risk. This seminal framework was later extended to convex risk measures by Föllmer and Schied [15] and Frittelli and Rosazza Gianin [16]. However, coherent and convex risk measures are static, meaning that they are not updated as new information becomes available. This motivates the development of dynamic risk measures (DRMs), including dynamic coherent risk measures (coherent DRMs) (see [2, 32]) and dynamic convex risk measures (convex DRMs) (see [17, 11, 23, 21, 3, 4, 14]), etc. Moreover, several other axiomatic notions in finance are closely related to DRMs. For instance, filtration-consistent nonlinear expectations ($\mathcal{F}$-expectations) introduced by Coquet et al. [8] have a one-to-one correspondence with a class of DRMs (which we call regular DRMs in the present paper). The dynamic concave utilities introduced by Delbaen et al. [10] have a one-to-one correspondence with convex DRMs.
A question raised in [8] can be generally described in the language of DRMs as follows: for a DRM $\rho$, can we find a function $g$ and a domain D{D} such that for all $\xi\in D$,
$\rho_{t}(\xi)={\cal{E}}^{g}[-\xi|{\cal{F}}_{t}],\quad\forall t\in[0,T]?$ (1.1)
Here, ${\cal{E}}^{g}[\cdot]$ is the $g$-expectation defined by the solution of the backward stochastic differential equation (BSDE) with generator $g$. This representation problem was considered theoretically very interesting and practically important by Peng in his ICM lecture (see [28, Page 403]). A positive answer means that any DRM satisfying certain conditions can be computed by solving a corresponding BSDE. This problem has been investigated for regular DRMs (i.e., for $\mathcal{F}$-expectations) in several settings. In [8], the authors obtained the first result, which shows that (1.1) holds for a Lipschitz continuous function $g$ on the domain $L^{2}(\mathcal{F}_{T})$ under the following domination condition:
$\rho_{t}(\xi)-\rho_{t}(\eta)\leqslant{\cal{E}}^{\mu,0}[\eta-\xi|{\cal{F}}_{t}],\quad\forall\xi,\eta\in L^{2}(\mathcal{F}_{T}),\ \forall t\in[0,T],$ (1.2)
[…]
Convex and coherent DRMs are two fundamental classes of risk measures. The main objective of this paper is to study the representation problem (1.1) for convex (resp. coherent) DRMs under the following domination condition:
${\cal{E}}^{-\mu,-\nu}[-\xi|{\cal{F}}_{t}]\leqslant\rho_{t}(\xi)\leqslant{\cal{E}}^{\mu,\nu}[-\xi|{\cal{F}}_{t}],\quad\forall\xi\in\mathcal{L}^{\exp}_{T}\ (\text{resp.}\ L^{2}(\mathcal{F}_{T})),\ \forall t\in[0,T],$ (1.3)
where ${\cal{E}}^{\mu,\nu}$ is the $g$-expectation defined by the BSDE with generator $\mu|z|+\nu|z|^{2},\mu\geqslant 0,\nu\geqslant 0$. Since $\rho_{t}(0)=0$ and ${\cal{E}}^{-\mu,0}[\cdot|{\cal{F}}_{t}]=-{\cal{E}}^{\mu,0}[-\cdot|{\cal{F}}_{t}]$, it follows that (1.2) implies ${\cal{E}}^{-\mu,0}[-\xi|{\cal{F}}_{t}]\leqslant\rho_{t}(\xi)\leqslant{\cal{E}}^{\mu,0}[-\xi|{\cal{F}}_{t}]$. Thus, (1.2) is a special case of (1.3). For convex (resp. coherent) DRMs, we prove that under (1.3), (1.1) holds on $\mathcal{L}^{\exp}_{T}$ (resp. $L^{2}(\mathcal{F}_{T})$) for a generator $g$ which is convex (resp. sublinear) in the variable $z$ and has a quadratic (resp. linear) growth (see Theorem 3.3 and Corollary 3.4). Compared to the works of [18] and [36], our result holds for unbounded terminal variables and does not require $\rho$ to have independent increments or satisfy additional convergence properties, and (1.3) seems to be easier to verify for convex DRMs in applications. Our proof follows the domination idea developed by [8] in the Lipschitz case. However, since some results that are crucial in the proof of the representation theorem in [8] generally fail to hold in non-Lipschitz settings, applying the domination argument in such cases, particularly the quadratic growth case, is really not easy. New domination conditions need to be found, and new strategies and techniques have to be developed, as seen in [18, 38, 36]. Inspired by the well-posedness results for quadratic BSDEs in Briand and Hu [5] and Fan et al. [12], we use the convexity property to obtain a $\theta$-domination for convex DRMs. This is then used to derive a crucial convergence property of convex DRMs. Combining this convergence, an exponential transformation and localization techniques, we establish a Doob-Meyer decomposition for a $\rho$-supermartingale with the special structure: $Y_{t}+z\cdot B_{t}$, where $Y_{t}$ is bounded. This decomposition, together with the aforementioned convergence result, enables us to complete the proof.
A dual representation of a convex DRM $\rho$ has the form:
$\rho_{t}(\xi)=\operatorname{ess\,sup}_{Q\in\mathcal{Q}}\left(E_{Q}[-\xi|\mathcal{F}_{t}]+\alpha(Q)\right),\quad\forall t\in[0,T],$ (1.4)
where $\xi\in L^{0}(\mathcal{F}_{T})$, $\mathcal{Q}$ is a set of probability measures and $\alpha$ is a penalty function. A dual representation of a coherent DRM $\rho$ has the form:
$\rho_{t}(\xi)=\operatorname{ess\,sup}_{Q\in\mathcal{Q}_{1}}E_{Q}[-\xi|\mathcal{F}_{t}],\quad\forall t\in[0,T],$ (1.5)
where $\xi\in L^{0}(\mathcal{F}_{T})$ and $\mathcal{Q}_{1}$ is a set of probability measures. For further details on the dual representation of convex DRMs, we refer to Detlefsen and Scandolo [11], Bino-Nadal [4], Föllmer and Schied [14], and the references therein. [10] provided a representation for the penalty function $\alpha$ in (1.4). They proved that under some additional assumptions, there exists a suitable nonnegative function $f$ such that for all $\xi\in L^{\infty}(\mathcal{F}_{T})$,
$\rho_{t}(\xi)=\operatorname{ess\,sup}_{Q^{q}\in\mathcal{Q}}E_{Q^{q}}\left[-\xi-\int_{t}^{T}f(s,q_{s})ds|\mathcal{F}_{t}\right],\quad\forall t\in[0,T],$ (1.6)
where $\mathcal{Q}:=\{Q^{q}:Q^{q}$ is a probability equivalent to $P$ such that $dQ/dP=\exp\{\int_{0}^{t}q_{s}ds-\frac{1}{2}\int_{0}^{t}|q_{s}|^{2}ds\}\}.$ Using the representation theorem established in this paper (see (3.2) in Theorem 3.3) and some results of Fan et al. [13], we show that any convex DRM satisfying (1.3) admits a dual representation as in (1.6). Compared to [10], our result holds for unbounded terminal variables, and the function $f$ can be determined by the Legendre-Fenchel transformation of the generator $g$ in (3.2). Furthermore, for coherent DRMs satisfying (1.3), we obtain a dual representation as in (1.5), where the terminal variable $\xi$ is unbounded and $\mathcal{Q}$ is characterized by the subdifferential $\partial g(t,0)$. Our method relies on the representation (3.2) and therefore requires (1.3).
The paper is organized as follows. In Section 2, we show some results on quadratic $g$-expectations. In Section 3, we introduce the main results of this paper. In Section 4, we prove some properties of convex DRMs. In Section 5, we establish a nonlinear Doob-Meyer decomposition. In Section 6, we provide the proofs of the main results. The Appendix contains an auxiliary proof.
2 Quadratic $g$-expectations
[…]
$L^{0}({\cal{F}}_{t}):=\{\xi:\ {\cal{F}}_{t}$-measurable random variable}\}, $t\in[0,T]$;
$L^{r}({\mathcal{F}}_{t}):=\{\xi\in L^{0}({\cal{F}}_{t}):$ ${{E}}\left[|\xi|^{r}\right]<\infty\},\ r\geqslant 1,\ t\in[0,T]$;
$L^{\infty}({\mathcal{F}}_{t}):=\{\xi\in L^{0}({\cal{F}}_{t}):$ $\|\xi\|_{\infty}:=\operatorname{ess\,sup}_{\omega\in\Omega}|\xi(\omega)|<\infty\},\ t\in[0,T]$;
${\mathcal{C}}:=\{(\psi_{t})_{t\in[0,T]}:$ continuous and $({\cal{F}}_{t})$-adapted process}\};
${\mathcal{S}}^{r}:=\{(\psi_{t})_{t\in[0,T]}\in{\mathcal{C}}:$ $E[{\mathrm{sup}}_{0\leqslant t\leqslant T}|\psi_{t}|^{r}]<\infty\},\ r\geqslant 1;$
${\mathcal{S}}^{\infty}:=\{(\psi_{t})_{t\in[0,T]}\in{\mathcal{C}}:$ $\|\psi_{t}\|_{\mathcal{S}^{\infty}}:=\operatorname{ess\,sup}_{(\omega,t)\in\Omega\times[0,T]}|\psi_{t}(\omega)|<\infty\}$;
$\mathcal{L}^{\exp}_{t}:=\{\xi:\exp(|\xi|)\in\bigcap_{r>1}L^{r}({\mathcal{F}}_{t})\},\ t\in[0,T]$;
$\mathcal{L}^{\exp}_{\mathcal{F}}:=\{(\psi_{t})_{t\in[0,T]}:\exp(|\psi_{t}|)\in\bigcap_{r>1}\mathcal{S}^{r}\}$.
${\mathcal{A}}:=\{(\psi_{t})_{t\in[0,T]}:$ increasing, RCLL and $({\cal{F}}_{t})$-adapted R-valued process with $\psi_{0}=0\}$;
${H}_{d}^{r}:=\{(\psi_{t})_{t\in[0,T]}:$ $\textbf{R}^{d}$-valued, $({\cal{F}}_{t})$-progressively measurable and $\int_{0}^{T}|\psi_{t}|^{r}dt<\infty\},\ r\geqslant 1;$
${\cal{H}}_{d}^{r}:=\{(\psi_{t})_{t\in[0,T]}\in{H}^{2}:$ $E[(\int_{0}^{T}|\psi_{t}|^{2}dt)^{\frac{r}{2}}]<\infty\},\ r\geqslant 1;$
${\cal{H}}_{d}^{BMO}:=\{(\psi_{t})_{t\in[0,T]}\in{\mathcal{H}}_{d}^{2}:$ $\|\psi_{t}\|_{BMO}:=\sup_{\tau\in{\cal{T}}_{0,T}}\|E[\int_{\tau}^{T}|\psi_{t}|^{2}dt|{\cal{F}}_{\tau}]\|_{\infty}^{1/2}<\infty\};$
${\cal{G}}:=\{g(\omega,t,z):\Omega\times[0,T]\times\mathbf{R}^{\mathit{d}}\longmapsto\mathbf{R}$ and is measurable with respect to ${\cal{P}}\otimes{\cal{B}}(\mathbf{R}^{d})\};$
${\cal{G}}^{\mu,\nu}:=\{g\in{\cal{G}}:$ $dt\times dP$-$a.e.,$ $\forall z\in\mathbf{R}^{\mathit{d}},\ |g(t,z)|\leqslant\mu|z|+\nu|z|^{2}\};$
${\cal{G}}^{\mu,\nu}_{\theta}:=\{g\in{\cal{G}}:$ $dt\times dP$-$a.e.,$ $\forall z,\tilde{z}\in\mathbf{R}^{d}$, $\forall\theta\in(0,1)$, $g(t,z)-\theta g(t,\tilde{z})\leqslant\mu|z-\tilde{z}|+\frac{\nu}{1-\theta}|z-\tilde{z}|^{2}\}$;
${\cal{G}}_{conv}:=\{g\in{\cal{G}}:$ $dt\times dP$-$a.e.,$ $g(t,z)$ is convex in z};z\};
${\cal{G}}_{subl}:=\{g\in{\cal{G}}:$ $dt\times dP$-$a.e.,$ $\forall z\in\mathbf{R}^{d},\forall\beta\geqslant 0$, $g(t,\beta z)=\beta g(t,z)$, and $\forall z_{1},z_{2}\in\mathbf{R}^{d},g(t,z_{1}+z_{2})\leqslant g(t,z_{1})+g(t,z_{2})\}.$
Remark 2.1.
For $g\in{\cal{G}}^{\mu,\nu}\cap{\cal{G}}^{\mu,\nu}_{\theta}$, we have $dt\times dP$-$a.e.,$ for all $z,\tilde{z}\in\mathbf{R}^{d}$ and $\theta\in(0,1)$,
[…]
$\leqslant\mu|z-\tilde{z}|+\frac{\nu}{1-\theta}|z-\tilde{z}|^{2}+(1-\theta)(\mu+\nu)(1+|z|^{2}+|\tilde{z}|^{2}).$
which implies that
$\lim_{\tilde{z}\rightarrow z}|g(t,z)-g(t,\tilde{z})|\leqslant(1-\theta)(\mu+\nu)(1+2|z|^{2})\rightarrow 0,\quad\text{as}\ \theta\rightarrow 1.$
It follows that for all $g\in{\cal{G}}^{\mu,\nu}\cap{\cal{G}}^{\mu,\nu}_{\theta}$, $dt\times dP$-$a.e.,$ $g(t,z)$ is continuous in $z$. Furthermore, from [12, Proposition 2.3(i)], it follows that ${\cal{G}}^{\mu,\nu}\cap{\cal{G}}_{conv}\subset{\cal{G}}^{\mu,\nu}\cap{\cal{G}}^{\mu,\nu}_{\theta}.$
[…]
$Y_{t}=\xi+\int_{t}^{T}g\left(s,Z_{s}\right)ds-\int_{t}^{T}Z_{s}\cdot dB_{s},\ \ t\in[0,T].$
[…]
Definition 2.2.
[…]
Since ${\cal{G}}^{\mu,\nu}\cap{\cal{G}}_{conv}\subset{\cal{G}}^{\mu,\nu}\cap{\cal{G}}^{\mu,\nu}_{\theta}$ (see Remark 2.1), it is clear that the two cases of Definition 2.2 are consistent. In the following, we list some properties of quadratic $g$-expectations. For more results on $g$-expectations, we refer to [27, 29, 30, 34, 33, 20, 25, 19, 35, 36] and the references therein.
Lemma 2.3.
Let $\xi\in\mathcal{L}^{\exp}_{T}$ and $g,f\in{\cal{G}}^{\mu,\nu}\cap{\cal{G}}^{\mu,\nu}_{\theta}$. Then the following hold:
(i) For all $t\in[0,T]$ and $s\in[0,t]$, ${\cal{E}}^{g}[{\cal{E}}^{g}[\xi|{\cal{F}}_{t}]|{\cal{F}}_{s}]={\cal{E}}^{g}[\xi|{\cal{F}}_{s}]$;
(ii) For all $t\in[0,T]$ and $\zeta\in\mathcal{L}^{\exp}_{t}$, ${\cal{E}}^{g}[\xi+\zeta|{\cal{F}}_{t}]={\cal{E}}^{g}[\xi|{\cal{F}}_{t}]+\zeta$;
(iii) For all $t\in[0,T]$ and $\zeta\in\mathcal{L}^{\exp}_{t}$, ${\cal{E}}^{g}[\zeta|{\cal{F}}_{t}]=\zeta$;
(iv) For $t\in[0,T]$, set $\tau_{t}:=\{s\geqslant 0:|B_{t+s}-B_{t}|>1\}\wedge(T-t)$. Then, for all $z\in\mathbf{R}^{d}$,
$\lim_{\varepsilon\rightarrow 0^{+}}{\cal{E}}^{g}[z\cdot(B_{t+\varepsilon\wedge\tau_{t}}-B_{t})|{\cal{F}}_{t}]=g(t,z),\ \ dt\text{-}a.e.;$
(v) For all $\beta>1$, there exists a constant $C>0$ depending only on $\beta$, $\mu$, $\nu$ and $T$, such that
$E\left[\exp\left(2(\mu+\nu)\beta\sup_{t\in[0,T]}|{\cal{E}}^{g}[\xi|{\cal{F}}_{t}]|\right)\right]\leqslant CE\left[\exp\left(2(\mu+\nu)\beta|\xi|\right)\right].$
(vi) For all $\eta\in\mathcal{L}^{\exp}_{T}$, if $\xi\leqslant\eta$, then for all $t\in[0,T]$, ${\cal{E}}^{g}[\xi|{\cal{F}}_{t}]\leqslant{\cal{E}}^{g}[\eta|{\cal{F}}_{t}]$;
(vii) $g(t,\cdot)\leqslant f(t,\cdot)$, $dt\times dP$-a.e., if and only if for all $\eta\in\mathcal{L}^{\exp}_{T}$,
${\cal{E}}^{g}[\eta|{\cal{F}}_{t}]\leqslant{\cal{E}}^{f}[\eta|{\cal{F}}_{t}],\quad\forall t\in[0,T];$
(viii) $g\in\mathcal{G}_{conv}$ if and only if for all $\eta,\zeta\in\mathcal{L}^{\exp}_{T}$ and $\theta\in(0,1)$,
${\cal{E}}^{g}[\theta\eta+(1-\theta)\zeta|{\cal{F}}_{t}]\leqslant\theta{\cal{E}}^{g}[\eta|{\cal{F}}_{t}]+(1-\theta){\cal{E}}^{g}[\zeta|{\cal{F}}_{t}],\quad\forall t\in[0,T];$
(ix) $g\in\mathcal{G}_{subl}$ if and only if for all $\eta,\zeta\in\mathcal{L}^{\exp}_{T}$ and $\beta\geqslant 0$,
${\cal{E}}^{g}[\beta\eta|{\cal{F}}_{t}]=\beta{\cal{E}}^{g}[\eta|{\cal{F}}_{t}]\ \ \text{and}\ \ {\cal{E}}^{g}[\eta+\zeta|{\cal{F}}_{t}]\leqslant{\cal{E}}^{g}[\eta|{\cal{F}}_{t}]+{\cal{E}}^{g}[\zeta|{\cal{F}}_{t}],\quad\forall t\in[0,T].$
Proof.
By the uniqueness of solutions, we obtain (i) and (ii). Since for all $t\in[0,T]$, ${\cal{E}}^{g}[0|{\cal{F}}_{t}]=0$, by (ii), we obtain (iii). (iv) is from Zheng [37, Corollary 3.2]. Since $|g(t,z)|\leqslant(\mu+\nu)(1+|z|^{2})$, (v) can be derived from [5, Corollary 4]. By the uniqueness of solutions and the comparison theorem in [36, Theorem 2.7(i)], we get (vi). In view of (iv) and the comparison theorem in [36, Theorem 2.7(ii)], by similar arguments as in [36, Proposition 3.5], we can get (vii)-(ix). ∎
Definition 2.4.
Let $g\in{\cal{G}}^{\mu,\nu}\cap{\cal{G}}^{\mu,\nu}_{\theta}$. A process $(Y_{t})_{t\in[0,T]}$ with $Y_{t}\in\mathcal{L}^{\exp}_{t}$ for $t\in[0,T],$ is called a $g$-martingale (resp. $g$-supermartingale, $g$-submartingale), if for any $0\leqslant s\leqslant t\leqslant T,$ we have ${\cal{E}}^{g}[Y_{t}|{\cal{F}}_{s}]=Y_{s}$, (resp. $\leqslant,\ \geqslant$).
For convenience, we denote $g^{\mu,\nu}(z):=\mu|z|+\nu|z|^{2}$, $g^{-\mu,-\nu}(z):=-\mu|z|-\nu|z|^{2}$, $\mathcal{E}^{\mu,\nu}:=\mathcal{E}^{g^{\mu,\nu}}$ and $\mathcal{E}^{-\mu,-\nu}:=\mathcal{E}^{g^{-\mu,-\nu}}$. The path regularity of bounded quadratic $g$-submartingales was obtained by Ma and Yao [25]. The following is a path regularity result of unbounded $g^{\mu,\nu}$-submartingales.
Lemma 2.5.
Let $Y_{t}$ be a $g^{\mu,\nu}$-submartingale (resp. $g^{-\mu,-\nu}$-supermartingale). Then $Y_{t}$ admits an RCLL modification.
Proof.
Let $Y_{t}$ be a $g^{\mu,\nu}$-submartingale. For any $0\leqslant s\leqslant t\leqslant T$, by Zheng et al. [39, Corollary 3.6], we have
$Y_{s}\leqslant\mathcal{E}^{\mu,\nu}[Y_{t}|{\cal{F}}_{s}]=\frac{1}{2\nu}\ln(E_{Q}[\exp(2\nu Y_{t})|{\cal{F}}_{s}]).$
Here, $Q$ is a probability measure such that $dQ/dP=\exp\{\int_{0}^{T}b_{s}\cdot dB_{s}-\frac{1}{2}\int_{0}^{T}|b_{s}|^{2}ds\}$, where $b_{t}$ is a progressively measurable process such that $|b_{t}|\leqslant\mu.$ This implies that for any $0\leqslant s\leqslant t\leqslant T$, $\exp(2\nu Y_{s})\leqslant E_{Q}[\exp(2\nu Y_{t})|{\cal{F}}_{s}]$. Thus, $\exp(2\nu Y_{t})$ is a classical submartingale under $Q$. Then by the classical martingale theory, we get that $Y_{t}$ admits an RCLL modification.
Let $Y_{t}$ be a $g^{-\mu,-\nu}$-supermartingale. Since for any $0\leqslant s\leqslant t\leqslant T$,
$Y_{s}\leqslant-{\cal{E}}^{-\mu,-\nu}[Y_{t}|{\cal{F}}_{s}]={\cal{E}}^{\mu,\nu}[-Y_{t}|{\cal{F}}_{s}],$
we get that $Y_{t}$ is a $g^{\mu,\nu}$-submartingale. Thus, $Y_{t}$ admits an RCLL modification. ∎
Remark 2.6.
In view of Lemma 2.5, for any $g^{\mu,\nu}$-submartingale (resp. $g^{-\mu,-\nu}$-supermartingale), we will always take its RCLL version.
At the end of this section, we give a Doob-Meyer decomposition of $g^{\mu,\nu}$-submartingales.
Lemma 2.7.
Let $Y_{t}$ be a $g^{\mu,\nu}$-submartingale such that $\sup_{t\in[0,T]}\exp(|Y_{t}|)\in\bigcap_{r>1}L^{r}(\mathcal{F}_{T})$. Then there exists a pair $(Z_{t},A_{t})\in H_{d}^{2}\times{\cal{A}}$, such that
$Y_{t}=Y_{T}+\int_{t}^{T}(\mu|Z_{s}|+\nu|Z_{s}|^{2})ds-A_{T}+A_{t}-\int_{t}^{T}Z_{s}\cdot dB_{s},\ \ t\in[0,T].$ (2.1)
Proof.
For $t\in[0,T]$, applying Itô’s formula to $\exp(2\nu\mathcal{E}^{\mu,\nu}[Y_{t}|{\cal{F}}_{s}])$ for $s\in[0,t]$, we have
$\exp(2\nu\mathcal{E}^{\mu,\nu}[Y_{t}|{\cal{F}}_{s}])=\mathcal{E}^{\mu,0}[\exp(2\nu Y_{t})|{\cal{F}}_{s}],\quad\forall s\in[0,t].$ (2.2)
Since for any $0\leqslant s\leqslant t\leqslant T$, $\exp(2\nu Y_{s})\leqslant\exp(2\nu\mathcal{E}^{\mu,\nu}[Y_{t}|{\cal{F}}_{s}])$, by (2.2), we get that $\exp(2\nu Y_{t})$ is a $g^{\mu,0}$-submartingale. Then, by Peng [29, Theorem 3.3], there exists a pair $(z_{t},a_{t})\in H_{d}^{2}\times{\cal{A}}$, such that
$\exp(2\nu Y_{t})=\exp(2\nu Y_{T})+\int_{t}^{T}\mu|z_{s}|ds-a_{T}+a_{t}-\int_{t}^{T}z_{s}\cdot dB_{s},\quad t\in[0,T].\\$
Let $a^{c}_{t}$ be the continuous part of $a_{t}$. Set $y_{t}:=\exp(2\nu Y_{t}),\ \Delta Y_{t}:=Y_{t}-Y_{t-},\ \Delta y_{t}:=y_{t}-y_{t-},\ \Delta a_{t}:=a_{t}-a_{t-},\ t\in(0,T]$. Applying Itô’s formula to $\frac{1}{2\nu}\ln(y_{t})$, we have
$Y_{t}=Y_{T}+$ $\int_{t+}^{T}\frac{1}{2\nu y_{s-}}\left(\mu|z_{s}|ds-da_{t}-z_{s}\cdot dB_{s}\right)+\int_{t+}^{T}\frac{|z_{s}|^{2}}{4\nu|y_{s}|^{2}}ds-\sum_{s\in(t,T]}\left(\Delta Y_{s}-\frac{1}{2\nu y_{s-}}\Delta y_{s}\right)$
$=Y_{T}+$ $\int_{t}^{T}\left(\mu\left|\frac{z_{s}}{2\nu y_{s}}\right|+\nu\left|\frac{z_{s}}{2\nu y_{s}}\right|^{2}\right)ds-\int_{t}^{T}\frac{z_{s}}{2\nu y_{s}}\cdot dB_{s}$
$-\int_{t}^{T}\frac{1}{2\nu y_{s}}da^{c}_{s}-\sum_{s\in(t,T]}\left(\frac{1}{2\nu y_{s-}}\Delta a_{s}+\Delta Y_{s}-\frac{1}{2\nu y_{s-}}\Delta y_{s}\right),\quad t\in[0,T].$
Since $\Delta y_{t}=\Delta a_{t}\geqslant 0$, by setting $Z_{t}:=\frac{z_{s}}{2\nu y_{s}}$ and $A_{t}:=\int_{0}^{t}\frac{1}{2\nu y_{s}}da^{c}_{s}+\sum_{s\in(0,t]}\Delta Y_{s}$, we get that $(Z_{t},A_{t})\in H_{d}^{2}\times{\cal{A}}$ and (2.1) holds. The proof is complete. ∎
3 Representations of dynamic convex risk measures
We first introduce some notions of dynamic risk measures (DRMs).
Definition 3.1.
A system of operators:
$\rho_{t}(\cdot):\ L^{0}(\mathcal{F}_{T})\longrightarrow L^{0}(\mathcal{F}_{t}),\quad t\in[0,T],$
is called a DRM on $L^{0}(\mathcal{F}_{T})$, if for all $\xi,\eta\in L^{0}(\mathcal{F}_{T})$, it satisfies (r1), (r2) and (r3):
(r1) Monotonicity: If $\xi\geqslant\eta$, then for all $t\in[0,T]$, $\rho_{t}(\xi)\leqslant\rho_{t}(\eta)$;
(r2) Time-consistency: For all $t\in[0,T]$ and $s\in[0,t]$, $\rho_{s}(-\rho_{t}(\xi))=\rho_{s}(\xi)$;
(r3) Constant preservation: If for some $t\in[0,T]$, $\xi\in L^{0}(\mathcal{F}_{t}),$ then $\rho_{t}(\xi)=-\xi$.
A DRM is called a convex DRM, if for all $\xi,\eta\in L^{0}(\mathcal{F}_{T})$, it satisfies (r4):
(r4) Convexity: For all $t\in[0,T]$ and $\theta\in(0,1)$, $\rho_{t}(\theta\xi+(1-\theta)\eta)\leqslant\theta\rho_{t}(\xi)+(1-\theta)\rho_{t}(\eta).$
A DRM is called a coherent DRM, if for all $\xi,\eta\in L^{0}(\mathcal{F}_{T})$, it satisfies (r5) and (r6):
(r5) Subadditivity: for all $t\in[0,T]$, $\rho_{t}(\xi+\eta)\leqslant\rho_{t}(\xi)+\rho_{t}(\eta)$;
(r6) Positive homogeneity: for all $\lambda\geqslant 0$ and $t\in[0,T]$, $\rho_{t}(\lambda\xi)=\lambda\rho_{t}(\xi)$.
A DRM is called a regular DRM, if for all $\xi\in L^{0}(\mathcal{F}_{T})$, it satisfies (r7):
(r7) Regular: for all $t\in[0,T]$ and $A\in{\mathcal{F}}_{t},$ $\rho_{t}(1_{A}\xi)=1_{A}\rho_{t}(\xi)$.
It can be seen that a coherent DRM is a convex DRM. From Lemma 2.3(viii), it is clear that if $g\in{\cal{G}}^{\mu,\nu}\cap{\cal{G}}_{conv}$, then ${\cal{E}}^{g}[-\cdot|{\cal{F}}_{t}]$ is a convex DRM on $\mathcal{L}^{\exp}_{T}$; From Rosazza Gianin [33], it follows that if $g\in{\cal{G}}^{\mu,0}\cap{\cal{G}}_{conv}$, then ${\cal{E}}^{g}[-\cdot|{\cal{F}}_{t}]$ is a convex DRM on $L^{2}(\mathcal{F}_{T})$.
Remark 3.2.
We present the correspondences between DRMs and several important notions in financial mathematics:
(i) If $\rho$ is a coherent DRM, then $\rho_{0}$ is a coherent risk measure (see [1]);
(ii) If $\rho$ is a convex DRM, then $\rho_{0}$ is a convex risk measure (see [15, 16]);
(iii) $\rho$ is a regular DRM if and only if $\rho(-\cdot)$ is an $\mathcal{F}$-expectation (see [8, 30]);
(iv) $\rho$ is a convex DRM if and only if $\rho$ is a dynamic concave utility (see [10]).
To solve Problem (1.1) for convex DRMs, we make the following assumption:
  • (A) For all $\xi\in\mathcal{L}^{\exp}_{T}$,
    ${\cal{E}}^{-\mu,-\nu}[-\xi|{\cal{F}}_{t}]\leqslant\rho_{t}(\xi)\leqslant{\cal{E}}^{\mu,\nu}[-\xi|{\cal{F}}_{t}],\quad\forall t\in[0,T].$ (3.1)
We have the following representation theorems:
Theorem 3.3.
Let $\rho$ be a convex DRM satisfying (A). Then there exists a unique $g\in{\cal{G}}^{\mu,\nu}\cap{\cal{G}}_{conv}$ such that for all $\xi\in\mathcal{L}^{\exp}_{T}$,
$\rho_{t}(\xi)={\cal{E}}^{g}[-\xi|{\cal{F}}_{t}],\quad\forall t\in[0,T].$ (3.2)
Moreover, the function $f(t,x):=\sup_{z\in\mathbf{R}^{d}}\{x\cdot z-g(t,z)\},\ (t,x)\in[0,T]\times\mathbf{R}^{d},$ is such that for all $\xi\in\mathcal{L}^{\exp}_{T}$,
$\rho_{t}(\xi)=\operatorname{ess\,sup}_{q\in\mathcal{Q}_{\xi,f}}E_{Q^{q}}\left[-\xi-\int_{t}^{T}f(s,q_{s})ds|\mathcal{F}_{t}\right],\quad\forall t\in[0,T],$ (3.3)
[…]
Corollary 3.4.
Let $\rho$ be a coherent DRM satisfying (A). Then there exists a unique $g\in{\cal{G}}^{\mu,0}\cap{\cal{G}}_{subl}$ such that for all $\xi\in\mathcal{L}^{\exp}_{T}$, (3.2) holds, and $g$ is also a unique function in ${\cal{G}}^{\mu,0}\cap{\cal{G}}_{subl}$ such that for all $\xi\in\mathcal{L}^{\exp}_{T}$,
$\rho_{t}(\xi)=\operatorname{ess\,sup}_{Q\in\mathcal{Q}_{g}}E_{Q}\left[-\xi|\mathcal{F}_{t}\right],\quad\forall t\in[0,T],$ (3.4)
where $\mathcal{Q}_{g}:=\{Q^{q}:$ $\frac{dQ^{q}}{dP}=\exp\{\int_{0}^{T}q_{s}\cdot dB_{s}-\frac{1}{2}\int_{0}^{T}|q_{s}|^{2}ds\}$ with $q_{t}\in H^{2}_{d}$ and $dt\times dP$-$a.e.,$ $q_{t}\in\partial g(t,0)\}$. Moreover, if for all $\xi\in L^{2}(\mathcal{F}_{T})$, $\rho$ satisfies (3.1) with $\nu=0$, then the function $g$ is such that for all $\xi\in L^{2}(\mathcal{F}_{T})$, both (3.2) and (3.4) hold.
The proofs of Theorem 3.3 and Corollary 3.4 will be given in Section 6. In the following, we make some remarks on Theorem 3.3 and Corollary 3.4.
Remark 3.5.
  • (i)
[…]
  • (ii)
    [10, Theorem 3.2] proved that any convex DRM admits a dual representation of the form (3.3) on $L^{\infty}({\cal{F}}_{T})$ under certain additional assumptions. Our Theorem 3.3, in contrast, provides a representation for unbounded terminal variables, where the function $f$ is explicitly given by the Legendre-Fenchel transformation of the generator $g$ in (3.2). Corollary 3.4 provides a dual representation for coherent DRMs in which the set of probability measures is characterized via the subdifferential $\partial g(t,0)$, a distinctive feature compared to those commonly found in the literature. Our method relies on (3.2) and therefore requires (A).
The following is a typical example of convex DRMs.
Example 3.6.
(Dynamic entropy risk measure) For $\xi\in\mathcal{L}^{\exp}_{T}$,
$\rho_{t}^{\mathrm{ent}}(\xi):=\frac{1}{2\nu}\ln E[\exp(-2\nu\xi)|\mathcal{F}_{t}],\quad t\in[0,T],$
is a convex DRM, called dynamic entropy risk measure. By Itô’s formula, we have
$\rho^{\mathrm{ent}}(\xi)=\mathcal{E}^{0,\nu}[-\xi|\mathcal{F}_{t}],\quad t\in[0,T],$
and thus $\rho$ satisfies (A). From Theorem 3.3, it follows that
$\rho^{\mathrm{ent}}_{t}(\xi)=\operatorname{ess\,sup}_{q\in\mathcal{Q}_{\xi,f}}E_{Q^{q}}\left[-\xi-\int_{t}^{T}\frac{|q_{s}|^{2}}{4\nu}ds|\mathcal{F}_{t}\right],\quad\forall t\in[0,T].$ (3.5)
In particular, when $\xi\in L^{\infty}(\mathcal{F}_{T})$, by (Appendix: The proof of (6.9)) in the Appendix, $\mathcal{Q}_{\xi,f}$ in (3.4) can be reduced to $\mathcal{Q}:=\{Q^{q}:$ $dQ^{q}/dP=\theta^{q}_{T}:=\int_{0}^{T}q_{s}\cdot dB_{s}-\frac{1}{2}\int_{0}^{T}|q_{s}|^{2}ds$ with $q_{t}\in\mathcal{H}_{d}^{BMO}\}.$ For $Q^{q}\in\mathcal{Q}$, the conditional relative entropy of $Q^{q}$ with respect to $P$:
$\mathbb{H}_{t}(Q^{q}|P):=E_{Q^{q}}\left[\ln\left(\frac{\theta^{q}_{T}}{\theta^{q}_{t}}\right)|\mathcal{F}_{t}\right]$ $=E_{Q^{q}}\left[\left(\int_{t}^{T}q_{s}\cdot dB_{s}-\frac{1}{2}\int_{t}^{T}|q_{s}|^{2}ds\right)|\mathcal{F}_{t}\right]$
$=E_{Q^{q}}\left[\left(\int_{t}^{T}q_{s}\cdot dB^{q}_{s}+\frac{1}{2}\int_{t}^{T}|q_{s}|^{2}ds\right)|\mathcal{F}_{t}\right]$
$=\frac{1}{2}E_{Q^{q}}\left[\int_{t}^{T}|q_{s}|^{2}ds|\mathcal{F}_{t}\right],\quad\forall t\in[0,T],$ (3.6)
where $B^{q}_{t}:=B_{t}-\int_{0}^{t}{q}_{s}ds$ is a standard Brownian motion under the probability $Q^{q}$. From (3.5) with $\mathcal{Q}_{\xi,f}$ replaced by $\mathcal{Q}$, and (3.6), it follows that for all $\xi\in L^{\infty}(\mathcal{F}_{T})$,
$\rho^{\mathrm{ent}}_{t}(\xi)=\operatorname{ess\,sup}_{Q^{q}\in\mathcal{Q}}\left(E_{Q^{q}}[-\xi|\mathcal{F}_{t}]-\frac{1}{2\nu}\mathbb{H}_{t}(Q^{q}|P)\right),\quad\forall t\in[0,T].$ (3.7)
A result similar to (3.7) can be found in [14].
4 Some properties of convex DRMs
In this section, we show some properties of convex DRMs.
Proposition 4.1.
Let $\rho$ be a convex DRM. Then the following hold:
(i) Translation invariance: For all $\xi\in L^{0}({\cal{F}}_{T})$, $t\in[0,T]$ and $\zeta\in L^{0}({\cal{F}}_{t})$, we have
$\rho_{t}(\xi+\zeta)=\rho_{t}(\xi)-\zeta;$
(ii) $\theta$-domination: For all $\xi,\eta\in L^{0}(\mathcal{F}_{T})$ and $\theta\in(0,1)$, we have
$\rho_{t}(\xi)-\theta\rho_{t}(\eta)\leqslant(1-\theta)\rho_{t}\left(\frac{\xi-\theta\eta}{1-\theta}\right),\quad\forall t\in[0,T];$
(iii) $L^{\infty}$-domination: For all $\xi,\eta\in L^{\infty}({\cal{F}}_{T})$ and $z\in\mathbf{R}^{d}$, we have
$\|\rho_{t}(\xi-z\cdot B_{T})-\rho_{t}(\eta-z\cdot B_{T})\|_{\mathcal{S}^{\infty}}\leqslant\|\xi-\eta\|_{\infty}.$
Proof.
(i) By (r4) and (r3), for all $\xi\in L^{0}({\cal{F}}_{T})$, $t\in[0,T]$ and $\zeta\in L^{0}({\cal{F}}_{t})$, we have
$\rho_{t}(\xi+\zeta)\leqslant\theta\rho_{t}\left(\frac{\xi}{\theta}\right)+(1-\theta)\rho_{t}\left(\frac{\zeta}{1-\theta}\right)=\theta\rho_{t}\left(\frac{\xi}{\theta}\right)-\zeta,\quad\forall\theta\in(0,1),$
and
$\rho_{t}(\xi)=\rho_{t}(\xi+\zeta-\zeta)\leqslant\theta\rho_{t}\left(\frac{\xi+\zeta}{\theta}\right)+(1-\theta)\rho_{t}\left(\frac{-\zeta}{1-\theta}\right)=\theta\rho_{t}\left(\frac{\xi+\zeta}{\theta}\right)+\zeta,\quad\forall\theta\in(0,1).$
By (r4), the function $h(x):=\rho_{t}(x\xi)$ is convex on $\mathbf{R}$, and thus continuous on $\mathbf{R}.$ Then, by letting $\theta\rightarrow 1$ in the two inequalities above, we obtain (i).
(ii) By (r4), for all $\xi,\eta\in L^{0}(\mathcal{F}_{T})$ and $\theta\in(0,1)$, we have
$\rho_{t}(\xi)-\theta\rho_{t}(\eta)$ $=\rho_{t}\left(\theta\eta+(1-\theta)\frac{\xi-\theta\eta}{1-\theta}\right)-\theta\rho_{t}(\eta)$
$\leqslant(1-\theta)\rho_{t}\left(\frac{\xi-\theta\eta}{1-\theta}\right),\quad\forall t\in[0,T].$
(iii) By (r1) and (i), for all $\xi,\eta\in L^{\infty}({\cal{F}}_{T})$, $t\in[0,T]$ and $z\in\mathbf{R}^{d}$, we have
$\rho_{t}(\xi-z\cdot B_{T})\leqslant\rho_{t}(\eta-z\cdot B_{T}-\|\xi-\eta\|_{\infty})=\rho_{t}(\eta-z\cdot B_{T})+\|\xi-\eta\|_{\infty}.$
By interchanging $\xi$ and $\eta$, we obtain (iii). ∎
Proposition 4.2.
Let $\rho$ be a convex DRM satisfying (A). Then the following hold:
(i) Convergence: For $\xi,\xi_{n}\in\mathcal{L}^{\exp}_{T}$ such that for all $\beta>1$, $\sup_{n\geqslant 1}E[\exp(\beta|\xi_{n}|))]<\infty$, if $\xi_{n}\rightarrow\xi,$ dPdP-$a.s.$, as $n\rightarrow\infty,$ then for all $\beta>1$,
$\lim_{n\rightarrow\infty}E\left[\exp\left(\sup_{t\in[0,T]}\beta|\rho_{t}(\xi_{n})-\rho_{t}(\xi)|\right)\right]=1;$
(ii) Regular: For all $\xi\in\mathcal{L}^{\exp}_{T}$, $t\in[0,T]$ and $A\in{\mathcal{F}}_{t}$, we have
$\rho_{t}(1_{A}\xi)=1_{A}\rho_{t}(\xi);$
(iii) Boundedness: For all $\xi\in L^{\infty}({\cal{F}}_{T})$ and $z\in\mathbf{R}^{d}$, we have
$\|\rho_{t}(\xi-z\cdot B_{T})-z\cdot B_{t}\|_{\mathcal{S}^{\infty}}<\infty.$
Proof.
(i) Since for all $\beta>1$, $\sup_{n\geqslant 1}E[\exp(\beta|\xi_{n}|))]<\infty$, by Lemma 2.3(v), we deduce that for all $\alpha>1$ and $\beta>1$,
$\sup_{n\geqslant 1}E\left[\exp\left(\beta\sup_{t\in{[0,T]}}{\cal{E}}^{\mu,\nu}[\alpha|\xi_{n}||{\cal{F}}_{t}]\right)\right]<\infty.$ (4.1)
Since $\xi_{n}\rightarrow\xi$, dPdP-$a.s.$, as $n\rightarrow\infty,$ by Fatou’s Lemma, we get that for all $\beta>1$,
$E[\exp(\beta|\xi|))]\leqslant\liminf_{n\rightarrow\infty}E[\exp(\beta|\xi_{n}|))]\leqslant\sup_{n\geqslant 1}E[\exp(\beta|\xi_{n}|))]<\infty,$ (4.2)
which together with Lemma 2.3(v) implies that for all $\alpha>1$ and $\beta>1$,
$E\left[\exp\left(\beta\sup_{t\in{[0,T]}}{\cal{E}}^{\mu,\nu}[\alpha|\xi||{\cal{F}}_{t}]\right)\right]<\infty.$ (4.3)
By Lemma 2.3(v), (4.2) and dominated convergence theorem, we get that for all $\alpha>1$ and $\beta>1$, there exists a constant $C>0$ depending only on $\beta$, $\mu$, $\nu$ and $T$, such that
$\lim_{n\rightarrow\infty}E\left[\exp\left(\beta\sup_{t\in{[0,T]}}{\cal{E}}^{\mu,\nu}\left[\alpha|\xi_{n}-\xi||{\cal{F}}_{t}\right]\right)\right]$ $\leqslant\lim_{n\rightarrow\infty}CE\left[\exp\left(\frac{\beta}{(\mu+\nu)\wedge 1}2(\mu+\nu)\alpha|\xi_{n}-\xi|\right)\right]$
$=C.$ (4.4)
By Lemma 2.3(vi) and (A), we have
$-{\cal{E}}^{\mu,\nu}[|\xi||{\cal{F}}_{t}]={\cal{E}}^{-\mu,-\nu}[-|\xi||{\cal{F}}_{t}]\leqslant{\cal{E}}^{-\mu,-\nu}[-\xi|{\cal{F}}_{t}]\leqslant\rho_{t}(\xi)\leqslant{\cal{E}}^{\mu,\nu}[-\xi|{\cal{F}}_{t}]\leqslant{\cal{E}}^{\mu,\nu}[|\xi||{\cal{F}}_{t}].$ (4.5)
By Proposition 4.1(ii), (4.5) and Lemma 2.3(viii)(vi), for all $n\geqslant 1$, $t\in[0,T]$, and $\theta\in(0,1)$, we have
$\rho_{t}(\xi_{n})-\rho_{t}(\xi)$ $\leqslant\rho_{t}(\xi_{n})-\theta\rho_{t}(\xi)-(1-\theta)\rho_{t}(\xi)$
$\leqslant(1-\theta)\rho_{t}\left(\frac{\xi_{n}-\theta\xi}{1-\theta}\right)+(1-\theta)|\rho_{t}(\xi)|$
$\leqslant(1-\theta){\cal{E}}^{\mu,\nu}\left[|\xi_{n}|+\frac{\theta}{1-\theta}|\xi_{n}-\xi||{\cal{F}}_{t}\right]+(1-\theta){\cal{E}}^{\mu,\nu}[|\xi||{\cal{F}}_{t}]$
$\leqslant\frac{1-\theta}{2}\left({\cal{E}}^{\mu,\nu}[2|\xi_{n}||{\cal{F}}_{t}]+{\cal{E}}^{\mu,\nu}\left[\frac{2\theta}{1-\theta}|\xi_{n}-\xi||{\cal{F}}_{t}\right]\right)+(1-\theta){\cal{E}}^{\mu,\nu}[|\xi||{\cal{F}}_{t}]$
$\leqslant(1-\theta)\left({\cal{E}}^{\mu,\nu}[2|\xi_{n}||{\cal{F}}_{t}]+{\cal{E}}^{\mu,\nu}[2|\xi||{\cal{F}}_{t}]+{\cal{E}}^{\mu,\nu}\left[\frac{2\theta}{1-\theta}|\xi_{n}-\xi||{\cal{F}}_{t}\right]\right).$
By interchanging $\xi_{n}$ and $\xi$, we have
$|\rho_{t}(\xi_{n})-\rho_{t}(\xi)|\leqslant(1-\theta)\left({\cal{E}}^{\mu,\nu}[2|\xi_{n}||{\cal{F}}_{t}]+{\cal{E}}^{\mu,\nu}[2|\xi||{\cal{F}}_{t}]+{\cal{E}}^{\mu,\nu}\left[\frac{2\theta}{1-\theta}|\xi_{n}-\xi||{\cal{F}}_{t}\right]\right).$
Then, by Jensen’s inequality and Hölder’s inequality, we get that for all $n\geqslant 1$ and $\theta\in(0,1)$,
$E\left[\exp\left(\sup_{t\in{[0,T]}}\beta|\rho_{t}(\xi_{n})-\rho_{t}(\xi)|\right)\right]$
$\leqslant E\left[\exp\left((1-\theta)\sup_{t\in{[0,T]}}\beta\left({\cal{E}}^{\mu,\nu}[2|\xi_{n}||{\cal{F}}_{t}]+{\cal{E}}^{\mu,\nu}[2|\xi||{\cal{F}}_{t}]+{\cal{E}}^{\mu,\nu}\left[\frac{2\theta}{1-\theta}|\xi_{n}-\xi||{\cal{F}}_{t}\right]\right)\right)\right]$
$\leqslant\left[E\exp\left(\sup_{t\in{[0,T]}}\beta\left({\cal{E}}^{\mu,\nu}[2|\xi_{n}||{\cal{F}}_{t}]+{\cal{E}}^{\mu,\nu}[2|\xi||{\cal{F}}_{t}]+{\cal{E}}^{\mu,\nu}\left[\frac{2\theta}{1-\theta}|\xi_{n}-\xi||{\cal{F}}_{t}\right]\right)\right)\right]^{(1-\theta)}$
$\leqslant\left[\sup_{n\geqslant 1}E\left[\exp\left(4\beta\sup_{t\in{[0,T]}}{\cal{E}}^{\mu,\nu}[2|\xi_{n}||{\cal{F}}_{t}]\right)\right]\right]^{\frac{(1-\theta)}{4}}\times\left[E\left[\exp\left(4\beta\sup_{t\in{[0,T]}}{\cal{E}}^{\mu,\nu}[2|\xi||{\cal{F}}_{t}]\right)\right]\right]^{\frac{(1-\theta)}{4}}$
$\ \ \ \ \times\left[E\left[\exp\left(2\beta\sup_{t\in{[0,T]}}{\cal{E}}^{\mu,\nu}\left[\frac{2\theta}{1-\theta}|\xi_{n}-\xi||{\cal{F}}_{t}\right]\right)\right]\right]^{\frac{(1-\theta)}{2}}$
In view of (4.1), (4.3) and (4), letting first $n\rightarrow\infty$ and then $\theta\rightarrow 1$, we obtain (i).
(ii) Set $\bar{\xi}_{n}:=1_{\{|\xi|\leqslant n\}}\xi,\ n\geqslant 1.$ As showed in Klöppel and Schweizer [23, Page 602], for all $t\in[0,T]$ and $A\in{\mathcal{F}}_{t}$, we have $1_{A}\rho_{t}(1_{A}\bar{\xi}_{n})=1_{A}\rho_{t}(\bar{\xi}_{n})$, which with the fact that $\rho_{t}(0)=0$, implies
$\rho_{t}(1_{A}\bar{\xi}_{n})=1_{A}\rho_{t}(1_{A}\bar{\xi}_{n})+1_{A^{c}}\rho_{t}(1_{A}\bar{\xi}_{n})=1_{A}\rho_{t}(\bar{\xi}_{n})+1_{A^{c}}\rho_{t}(0)=1_{A}\rho_{t}(\bar{\xi}_{n}).$ (4.6)
By (i), we deduce that for all $p>1$ and $A\in{\mathcal{F}}_{t}$, $\lim_{n\rightarrow\infty}E[\sup_{t\in[0,T]}|\rho_{t}(1_{A}\bar{\xi}_{n})-\rho_{t}(1_{A}\xi)|^{p}]=0.$ From this and (4.6), we get (ii).
[…]
${\cal{E}}^{\mu,\nu}[-\xi+z\cdot B_{T}|{\cal{F}}_{t}]=-\xi+z\cdot B_{T}+\int_{t}^{T}g^{\mu,\nu}(Z_{s})ds-\int_{t}^{T}Z_{s}\cdot dB_{s},\quad t\in[0,T].$
Then, we have
${\cal{E}}^{\mu,\nu}[-\xi+z\cdot B_{T}|{\cal{F}}_{t}]-z\cdot B_{t}=-\xi+\int_{t}^{T}g^{\mu,\nu}(Z_{s})ds-\int_{t}^{T}(Z_{s}-z)\cdot dB_{s},\quad t\in[0,T],$ (4.7)
[…]
$0\leqslant g^{\mu,\nu}(Z_{t})\leqslant g^{\mu,\nu}(|Z_{t}-z|+|z|),$
by (4.7) and the comparison theorem (see [5, Theorem 5] and its proof), we get
$E[-\xi|{\cal{F}}_{t}]\leqslant{\cal{E}}^{\mu,\nu}[-\xi+z\cdot B_{T}|{\cal{F}}_{t}]-z\cdot B_{t}\leqslant y^{1}_{t},\quad\forall t\in[0,T].$
which implies that ${\cal{E}}^{\mu,\nu}[-\xi+z\cdot B_{T}|{\cal{F}}_{t}]-z\cdot B_{t}\in{\cal{S}}^{\infty}.$ This also gives
${\cal{E}}^{-\mu,-\nu}[-\xi+z\cdot B_{T}|{\cal{F}}_{t}]-z\cdot B_{t}=-({\cal{E}}^{\mu,\nu}[\xi-z\cdot B_{T}|{\cal{F}}_{t}]+z\cdot B_{t})\in{\cal{S}}^{\infty}.$
Then, by (A), we obtain (iii). ∎
The following shows that under (A), for all $\xi\in\mathcal{L}^{\exp}_{T},$ $\rho_{t}(\xi)$ is an Itô process and thus all the paths of $\rho_{t}(\xi)$ are continuous.
Proposition 4.3.
Let $\rho$ be a convex DRM satisfying (A). Then for all $\xi\in\mathcal{L}^{\exp}_{T},$ there exists a pair $(g_{t}^{\xi},Z_{t}^{\xi})\in H_{1}^{1}\times H_{d}^{2}$ such that
$\rho_{t}(\xi)=-\xi+\int_{t}^{T}g_{s}^{\xi}ds-\int_{t}^{T}Z_{s}^{\xi}\cdot dB_{s},\quad t\in[0,T],$ (4.8)
with $|g_{t}^{\xi}|\leqslant g^{\mu,\nu}(Z_{t}^{\xi})$, $dt\times dP$-$a.e.$ Moreover, for all $\xi,\eta\in\mathcal{L}^{\exp}_{T}$ and $\theta\in(0,1)$, we have $dt\times dP$-$a.e.,$
$\frac{g_{t}^{\xi}-\theta g_{t}^{\eta}}{1-\theta}\leqslant g^{\mu,\nu}\left(\frac{Z_{t}^{\xi}-\theta Z_{t}^{\eta}}{1-\theta}\right).$
Proof.
For all $\xi\in\mathcal{L}^{\exp}_{T}$ and $0\leqslant s\leqslant t\leqslant T$, by (A), we have $\rho_{s}(\xi)=\rho_{s}(-\rho_{t}(\xi))\leqslant{\cal{E}}^{\mu,\nu}[\rho_{t}(\xi)|{\cal{F}}_{s}]$ and
$\rho_{s}(\xi)=-\rho_{s}(-\rho_{t}(\xi))\leqslant-{\cal{E}}^{-\mu,-\nu}[\rho_{t}(\xi)|{\cal{F}}_{s}]={\cal{E}}^{\mu,\nu}[-\rho_{t}(\xi)|{\cal{F}}_{s}],$
which imply that both $\rho_{t}(\xi)$ and $\rho_{t}(\xi)$ are $g^{\mu,\nu}$-submartingales. From Lemma 2.7, it follows that there exist $(Z^{1}_{t},A^{1}_{t})$ and $(Z^{2}_{t},A^{2}_{t})$ in $H_{d}^{2}\times\cal{A}$ such that
$\rho_{t}(\xi)=-\xi+\int_{t}^{T}g^{\mu,\nu}(Z_{s}^{1})ds-A^{1}_{T}+A^{1}_{t}-\int_{t}^{T}Z_{s}^{1}\cdot dB_{s},\quad t\in[0,T],$ (4.9)
and
$\rho_{t}(\xi)=-\xi-\int_{t}^{T}g^{\mu,\nu}(-Z_{s}^{2})ds+A^{2}_{T}-A^{2}_{t}-\int_{t}^{T}-Z_{s}^{2}\cdot dB_{s},\quad t\in[0,T].$ (4.10)
Comparing the martingale parts and the bounded variation parts of (4.9) and (4.10), we get that $dt\times dP$-$a.e.,$
$Z_{t}^{1}$ $=-Z_{t}^{2},$
$g^{\mu,\nu}(Z_{t}^{1})dt-dA^{1}_{t}$ $=-g^{\mu,\nu}(-Z_{t}^{2})dt+dA^{2}_{t},$
which imply that $dt\times dP$-$a.e.,$
$2g^{\mu,\nu}(Z_{t}^{1})dt=dA^{1}_{t}+dA^{2}_{t}.$
Since $A^{1}_{t}$ and $A^{2}_{t}$ are both increasing and continuous, by the Lebesgue decomposition for increasing functions, we can get that $A^{1}_{t}$ and $A^{2}_{t}$ are both absolutely continuous. Thus, we can find two nonnegative processes $a^{1}_{t},a^{2}_{t}\in H_{1}^{1}$ such that $dt\times dP$-$a.e.,$ $dA^{1}_{t}=a^{1}_{t}dt$ and $dA^{2}_{t}=a^{2}_{t}dt.$ This implies $dt\times dP$-$a.e.,$
$2g^{\mu,\nu}(|Z_{t}^{1}|)=a^{1}_{t}+a^{2}_{t}.$ (4.11)
Set $Z^{\xi}_{t}:=Z_{t}^{1}$ and $g^{\xi}_{t}:=g^{\mu,\nu}(Z_{t}^{\xi})-a^{1}_{t}.$ From (4.9) and (4.11), it follows that (4.8) holds and $dt\times dP$-$a.e.,$ $|g^{\xi}_{t}|\leqslant g^{\mu,\nu}(Z_{t}^{\xi})$.
Moreover, by Proposition 4.1(ii) and (A), for $\xi,\eta\in\mathcal{L}^{\exp}_{T}$ and $\theta\in(0,1)$, we have
$\frac{\rho_{s}(\xi)-\theta\rho_{s}(\eta)}{1-\theta}$ $=\frac{\rho_{s}(-\rho_{t}(\xi))-\theta\rho_{s}(-\rho_{t}(\eta))}{1-\theta}$
$\leqslant\rho_{s}\left(\frac{-\rho_{t}(\xi)-\theta(-\rho_{t}(\eta))}{1-\theta}\right)$
$={\cal{E}}^{\mu,\nu}\left[\frac{\rho_{t}(\xi)-\theta\rho_{t}(\eta)}{1-\theta}|{\cal{F}}_{s}\right],\quad\forall 0\leqslant s\leqslant t\leqslant T.$
This means that $\frac{\rho_{t}(\xi)-\theta\rho_{t}(\eta)}{1-\theta}$ is a $g^{\mu,\nu}$-submartingale. Then by Lemma 2.7 again, there exists a pair $(z_{t},A_{t})\in H_{d}^{2}\times\cal{A}$ such that
$\frac{\rho_{t}(\xi)-\theta\rho_{t}(\eta)}{1-\theta}=\frac{-\xi-\theta(-\eta)}{1-\theta}+\int_{t}^{T}g^{\mu,\nu}(z_{t})ds-A_{T}+A_{t}-\int_{t}^{T}z_{t}\cdot dB_{s},\quad t\in[0,T].$
Moreover, by (4.8), we have
$\frac{\rho_{t}(\xi)-\theta\rho_{t}(\eta)}{1-\theta}=\frac{-\xi-\theta(-\eta)}{1-\theta}+\int_{t}^{T}\left(\frac{g_{s}^{\xi}-\theta g_{s}^{\eta}}{1-\theta}\right)ds-\int_{t}^{T}\left(\frac{Z_{s}^{\xi}-\theta Z_{s}^{\eta}}{1-\theta}\right)\cdot dB_{s},\quad t\in[0,T].$
Comparing the martingale parts and the bounded variation parts of the two equations above, we get that $dt\times dP$-$a.e.,$
$\frac{g_{t}^{\xi}-\theta g_{t}^{\xi}}{1-\theta}\leqslant g^{\mu,\nu}(z_{t})\quad\textrm{with}\quad z_{t}=\frac{Z_{t}^{\xi}-\theta Z_{t}^{\eta}}{1-\theta}.$
The proof is complete. ∎
Remark 4.4.
Let $\rho$ be a convex DRM satisfying (A). For $\xi\in\mathcal{L}^{\exp}_{T}$ and $\tau\in{\cal{T}}_{0,T}$, we set $\rho_{\tau}(\xi)(\omega):=\rho_{\tau(\omega)}(\xi)(\omega).$ Then, it is not hard to check that on the domain $\mathcal{L}^{\exp}_{T}$, all the results in (r1)-(r7), Proposition 4.1(i)(ii), Proposition 4.2(ii) and (4.8) still hold, when $t$ therein are replaced by any stopping time $\tau\in{\cal{T}}_{0,T}.$ We only show the case of Proposition 4.1(i). For all $\tau\in{\cal{T}}_{0,T}$, we can find a decreasing sequence $\{\tau_{n}\}_{n\geqslant 1}\subset{\cal{T}}_{0,T}$, which all take values in some finite set such that $\tau_{n}\searrow\tau$ as $n\rightarrow\infty.$ We assume that $\tau_{n}$ take values: $t^{n}_{1},\cdots,t^{n}_{n_{k}}$, by Proposition 4.1(i), for all $n\geqslant 1$ and $\xi\in\mathcal{L}^{\exp}_{T},\zeta\in\mathcal{L}^{\exp}_{\tau}$, we have
$\rho_{\tau_{n}}(\xi+\zeta)=\sum_{i=1}^{n_{k}}1_{\{\tau_{n}=t^{n}_{i}\}}\rho_{t^{n}_{i}}(\xi+\zeta)=\sum_{i=1}^{n_{k}}1_{\{\tau_{n}=t^{n}_{i}\}}\rho_{t^{n}_{i}}(\xi)-\zeta=\rho_{\tau_{n}}(\xi)-\zeta.$
Since all the paths of $\rho_{t}(\xi)$ are continuous, letting $n\rightarrow\infty,$ we obtain $\rho_{\tau}(\xi+\zeta)=\rho_{\tau}(\xi)-\zeta$.
5 A Doob-Meyer decomposition for $\rho$-supermartingales
In this section, we will establish a decomposition of Doob-Meyer’s type for $\rho$-supermartingales, which plays a key role in the proof of Theorem 3.3. We first introduce the notion of $\rho$-martingales (resp. $\rho$-supermartingales, $\rho$-submartingales).
Definition 5.1.
Let $\rho$ be a convex DRM, and let $Y_{t}$ be a progressively measurable process such that for all $t\in[0,T],$ $Y_{t}\in\mathcal{L}^{\exp}_{t}$. If for all $0\leqslant s\leqslant t\leqslant T,$ we have $\rho_{s}(-Y_{t})=Y_{s}$ (resp. $\leqslant,\ \geqslant$), then the process $Y_{t}$ is called a $\rho$-martingale (resp. $\rho$-supermartingale, $\rho$-submartingale).
We have the following optional stopping theorem:
Proposition 5.2.
Let $\rho$ be a convex DRM satisfying (A). Let $Y_{t}$ be a $\rho$-supermartingale (resp. $\rho$-submartingale). Then, for all $\sigma,\tau\in{\cal{T}}_{0,T},$
$\rho_{\sigma}(-Y_{\tau})\leqslant Y_{\sigma\wedge\tau},\ (\textrm{resp.}\geqslant Y_{\sigma\wedge\tau}).$
Proof.
By Proposition 4.2(ii) and the proof of [8, Lemma 3.5], for all $\xi,\eta\in\mathcal{L}^{\exp}_{T}$, $t\in[0,T]$ and $A\in{\mathcal{F}}_{t},$ we have
$\rho_{t}(1_{A}\xi+1_{A^{c}}\eta)=1_{A}\rho_{t}(\xi)+1_{A^{c}}\rho_{t}(\eta).$ (5.1)
We only prove the case that $Y_{t}$ is a $\rho$-supermartingale. In this case, we have for all $0\leqslant s\leqslant t\leqslant T,$ $\mathcal{E}^{-\mu,-\nu}[Y_{t}]\leqslant\rho_{s}(-Y_{t})\leqslant Y_{s}$. This implies that $Y_{t}$ is a $g^{-\mu,-\nu}$-supermartingale. Thus, each path of $Y_{t}$ is right continuous (see Remark 2.6). In view of this, (5.1) and the convergence in Proposition 4.2(i), following the proof of [30, Theorem 7.4], we can obtain the desired result. ∎
We have the following well-posedness result:
Proposition 5.3.
Let $\rho$ be a convex DRM satisfying (A). Let $Y_{t}\in\mathcal{S}^{\infty}$, $z\in\mathbf{R}^{d}$, and $Y_{t}+z\cdot B_{t}$ be a $\rho$-supermartingale. Then, for any $n\geqslant 1$, the following backward equation:
$y_{t}^{n}=\rho_{t}\left(-Y_{T}-z\cdot(B_{T}-B_{t})-\int_{t}^{T}n(Y_{s}-y_{s}^{n})ds\right),\quad t\in[0,T],$ (5.2)
admits a unique solution $y_{t}^{n}\in{\cal{S}}^{\infty}$. Moreover, the following hold:
(i) For all $n\geqslant 1$ and $t\in[0,T]$, $y_{t}^{n}\leqslant y_{t}^{n+1}\leqslant Y_{t};$
(ii) For all $n\geqslant 1$, $y^{n}_{t}+z\cdot B_{t}$ is a $\rho$-supermartingale.
Proof.
Step 1. For $n\geqslant 1$ and $y_{t}\in{\mathcal{S}}^{\infty}$, we set
$\phi(y_{t}):=\rho_{t}\left(-Y_{T}-z\cdot B_{T}-\int_{0}^{T}n(Y_{s}-y_{s})ds\right)-z\cdot B_{t}-\int_{0}^{t}n(Y_{s}-y_{s})ds,\quad t\in[0,T].$
By Proposition 4.2(iii), for any $y_{t},\tilde{y}_{t}\in{\mathcal{S}}^{\infty}$, we have $\phi(y_{t}),\phi(\tilde{y}_{t})\in{\cal{S}}^{\infty}$ and by Proposition 4.1(iii),
$\|\phi(y_{t})-\phi(\tilde{y}_{t})\|_{\mathcal{S}^{\infty}}\leqslant 2nT\|y_{t}-\tilde{y}_{t}\|_{\mathcal{S}^{\infty}}.$
This implies that if $T\leqslant\frac{1}{4n}$, then $\phi(\cdot)$ is a strict contraction on $\mathcal{S}^{\infty}$. Thus, when $T\leqslant\frac{1}{4n}$, by Proposition 4.1(i), we get that the equation (5.2) admits a unique solution $y_{t}^{n}\in{\cal{S}}^{\infty}$.
When $T>\frac{1}{4n}$, we set $0=T_{1}<T_{2}<\cdots<T_{m}=T$ such that for any $1<k\leqslant m$, $T_{k}-T_{k-1}\leqslant\frac{1}{4n}$. Since for any $1<k\leqslant m$, the equation (5.2) has a unique bounded solution on $T_{k-1},T_{k}]$, using the “patching-up” method in the proof of [18, Proposition 4.4], we conclude that the equation (5.2) admits a unique solution $y_{t}^{n}\in{\cal{S}}^{\infty}$.
Step 2. For $n\geqslant 1$, suppose that there exists ${t_{0}}\in[0,T)$ such that $P(y_{t_{0}}^{n}>Y_{t_{0}})>0$, which implies that there exists $k\geqslant 1$ such that
$P\left(y_{t_{0}}^{n}>Y_{t_{0}}+\frac{1}{k}\right)>0.$ (5.3)
Set $A_{1}:=\{y_{t_{0}}^{n}>Y_{t_{0}}+\frac{1}{k}\}$ and $\tau:=\{t\geqslant{t_{0}}:y_{t}^{n}\leqslant Y_{t}\}\wedge T$. Since $y_{T}^{n}=Y_{T}$, we have
$1_{A_{1}}y_{\tau}^{n}=1_{A_{1}}Y_{\tau}\quad\text{and}\quad 1_{A_{1}}y_{t}^{n}\geqslant 1_{A_{1}}Y_{t},\ t\in[{t_{0}},\tau].$ (5.4)
From Proposition 4.1(i) and (5.2), it follows that
$y_{t}^{n}+z\cdot B_{t}+\int_{0}^{t}n(Y_{s}-y^{n}_{s})ds=\rho_{t}\left(-Y_{T}-z\cdot B_{T}-\int_{0}^{T}n(Y_{s}-y_{s}^{n})ds\right),\quad\forall t\in[0,T],$ (5.5)
which implies that $y_{t}^{n}+z\cdot B_{t}+\int_{0}^{t}n(Y_{s}-y^{n}_{s})ds$ is a $\rho$-martingale. Then by Proposition 5.2, Proposition 4.1(i), Proposition 4.2(ii), (r1) and (5.4), we have
$1_{A_{1}}y_{t_{0}}^{n}$ $=1_{A_{1}}\rho_{t_{0}}\left(-y_{\tau}^{n}-z\cdot(B_{\tau}-B_{t_{0}})-\int_{t_{0}}^{\tau}n(Y_{s}-y_{s}^{n})ds\right)$
$=1_{A_{1}}\rho_{t_{0}}\left(-1_{A_{1}}y_{\tau}^{n}-1_{A_{1}}z\cdot(B_{\tau}-B_{t_{0}})-\int_{t_{0}}^{\tau}n(1_{A_{1}}Y_{s}-1_{A_{1}}y_{s}^{n})ds\right)$
$\leqslant 1_{A_{1}}\rho_{t_{0}}\left(-1_{A_{1}}Y_{\tau}-1_{A_{1}}z\cdot(B_{\tau}-B_{t_{0}})\right)$
$=1_{A_{1}}\rho_{t_{0}}\left(-Y_{\tau}-z\cdot B_{\tau}\right)-1_{A_{1}}z\cdot B_{t_{0}}$
$\leqslant 1_{A_{1}}Y_{t_{0}},$
which contradicts (5.3). Thus, for all $n\geqslant 1$ and $t\in[0,T]$, $y_{t}^{n}\leqslant Y_{t}.$
Step 3. For $n\geqslant 1$, suppose that there exists ${t_{0}}\in[0,T)$ such that $P(y_{t_{0}}^{n}>y_{t_{0}}^{n+1})>0$, which implies that there exists $k\geqslant 1$ such that
$P\left(y_{t_{0}}^{n}>y_{t_{0}}^{n+1}+\frac{1}{k}\right)>0.$ (5.6)
Set $A_{2}:=\{y_{t_{0}}^{n}>y_{t_{0}}^{n+1}+\frac{1}{k}\}$ and $\sigma:=\{t\geqslant{t_{0}}:y_{t}^{n}\leqslant y_{t}^{n+1}\}\wedge T$. We then have
$1_{A_{2}}y_{\sigma}^{n}=1_{A_{2}}y_{\sigma}^{n+1}\quad\text{and}\quad 1_{A_{2}}y_{t}^{n}\geqslant 1_{A_{2}}y_{t}^{n+1},\ t\in[{t_{0}},\sigma].$ (5.7)
By Proposition 5.2, Proposition 4.1(i), Proposition 4.2(ii), (5.7), the consequence of Step 2, (r1), and (5.5), we have
$1_{A_{2}}y_{t_{0}}^{n}$ $=1_{A_{2}}\rho_{t_{0}}\left(-y_{\sigma}^{n}-z\cdot(B_{\sigma}-B_{t_{0}})-\int_{t_{0}}^{\sigma}n(Y_{s}-y_{s}^{n})ds\right)$
$=1_{A_{2}}\rho_{t_{0}}\left(-1_{A_{2}}y_{\sigma}^{n}-1_{A_{2}}z\cdot(B_{\sigma}-B_{t_{0}})-\int_{t_{0}}^{\sigma}n(1_{A_{2}}Y_{s}-1_{A_{2}}y_{s}^{n})ds\right)$
$\leqslant 1_{A_{2}}\rho_{t_{0}}\left(-1_{A_{2}}y_{\sigma}^{n+1}-1_{A_{2}}z\cdot(B_{\sigma}-B_{t_{0}})-\int_{t_{0}}^{\sigma}(n+1)(1_{A_{2}}Y_{s}-1_{A_{2}}y_{s}^{n+1})ds\right)$
$=1_{A_{2}}\rho_{t_{0}}\left(-y_{\sigma}^{n+1}-z\cdot(B_{\sigma}-B_{t_{0}})-\int_{t_{0}}^{\sigma}(n+1)(Y_{s}-y_{s}^{n+1})ds\right)$
$=1_{A_{2}}y_{t_{0}}^{n+1},$
which contradicts (5.6). Thus, for all $n\geqslant 1$ and $t\in[0,T]$, $y_{t}^{n}\leqslant y_{t}^{n+1}.$ This with Step 2 implies (i).
Step 4. By (5.5), Proposition 4.1(i), the consequence of Step 2, and (r1), we have
$y_{s}^{n}+z\cdot B_{s}$ $=\rho_{s}\left(-y_{t}^{n}-z\cdot B_{t}-\int_{s}^{t}n(Y_{r}-y_{r}^{n})dr\right)\geqslant\rho_{s}(-y_{t}^{n}-z\cdot B_{t}),\quad\forall 0\leqslant s\leqslant t\leqslant T,$
which implies that $y_{t}^{n}+z\cdot B_{t}$ is a $\rho$-supermartingale. Thus, (ii) holds. ∎
We have the following decomposition theorem for $\rho$-supermartingales.
Theorem 5.4.
Let $\rho$ be a convex DRM satisfying (A). Let $Y_{t}\in\mathcal{S}^{\infty}$, $z\in\mathbf{R}^{d}$, and $Y_{t}+z\cdot B_{t}$ be a $\rho$-supermartingale. Then there exists a process $A_{t}\in{\cal{A}}$ and an increasing sequence $\{\tau_{n}\}_{n\geqslant 1}\subset{\cal{T}}_{0,T}$ satisfying $\tau_{n}\nearrow T$ as $n\rightarrow\infty$, such that for all $n\geqslant 1$, $Y_{\tau_{n}\wedge t}+z\cdot B_{\tau_{n}\wedge t}+A_{\tau_{n}\wedge t}$ is a $\rho$-martingale. Moreover, if $A_{T}\in L^{\infty}({\cal{F}}_{T})$, then $Y_{t}+z\cdot B_{t}+A_{t}$ is a $\rho$-martingale.
Proof.
By Proposition 5.3 and Proposition 4.1(i), the following equation:
$y_{t}^{n}+z\cdot B_{t}+\int_{0}^{t}n(Y_{s}-y_{s}^{n})ds=\rho_{t}\left(-Y_{T}-z\cdot B_{T}-\int_{0}^{T}n(Y_{s}-y_{s}^{n})ds\right),\quad t\in[0,T],$ (5.8)
admits a unique solution $y_{t}^{n}\in{\cal{S}}^{\infty}$ such that for all $n\geqslant 1$ and $t\in[0,T]$, $y_{t}^{n}\leqslant y_{t}^{n+1}\leqslant Y_{t}.$ Thus, there exists a constant $C>0$ such that $dt\times dP$-$a.e.$,
$\sup_{n\geqslant 1}|y^{n}_{t}|<C.$ (5.9)
For $n\geqslant 1$, set
$A_{t}^{n}:=\int_{0}^{t}n(Y_{s}-y_{s}^{n})ds,\quad t\in[0,T].$ (5.10)
This with (5.8) implies that for all $t\in[0,T],$
$\rho_{t}(-Y_{T}-z\cdot B_{T}-A_{T}^{n})=y_{t}^{n}+z\cdot B_{t}+A_{t}^{n}.$ (5.11)
From Proposition 4.3, it follows that there exists a pair $(g^{n}_{s},Z^{n}_{s})\in H^{1}_{1}\times H_{d}^{2}$ such that
$y_{t}^{n}+z\cdot B_{t}=Y_{T}+z\cdot B_{T}+A_{T}^{n}-A_{t}^{n}+\int_{t}^{T}g^{n}_{s}ds-\int_{t}^{T}Z^{n}_{s}\cdot dB_{s},\quad t\in[0,T],$ (5.12)
with
$|g^{n}_{t}|\leqslant(\mu+\nu)(1+|Z^{n}_{t}|^{2}),\quad dt\times dP\text{-}a.e.,$ (5.13)
and for $m\geqslant 1$ and $\theta\in(0,1)$,
$\frac{g_{t}^{n}-\theta g_{t}^{m}}{1-\theta}\leqslant(\mu+\nu)\left(1+\left|\frac{Z^{n}_{t}-\theta Z^{m}_{t}}{1-\theta}\right|^{2}\right),\quad dt\times dP\text{-}a.e.$ (5.14)
Then we have
Proposition 5.5.
$\lim_{n\rightarrow\infty}E\left[\sup_{t\in[0,T]}|y^{n}_{t}-Y_{t}|\right]\rightarrow 0,$ (5.15)
$\lim_{n,m\rightarrow\infty}E\left[\int_{0}^{T}|Z^{n}_{t}-Z^{m}_{t}|^{2}dt\right]\rightarrow 0,\quad and\quad\lim_{n,m\rightarrow\infty}E\left[\int_{0}^{T}|g^{n}_{t}-g^{m}_{t}|dt\right]\rightarrow 0.$ (5.16)
Proof.
For $n\geqslant 1$, $k>0$ and $\delta\in\mathcal{T}_{0,T},$ we set
${\tau^{k}_{\delta}}:=\inf\left\{t\geqslant\delta:\int_{\delta}^{t}|Z^{n}_{s}-z|^{2}ds\geqslant k\right\}\wedge T.$
For a constant $\beta<0$, applying Itô’s formula to $e^{\beta y^{n}_{t}}$ and by (5.13), we have
$e^{\beta y^{n}_{\delta}}$ $+\frac{1}{2}\beta^{2}\int_{\delta}^{\tau^{k}_{\delta}}e^{\beta y^{n}_{s}}|Z^{n}_{s}-z|^{2}ds-\beta\int_{\delta}^{\tau^{k}_{\delta}}e^{\beta y^{n}_{s}}dA_{s}$
$\leqslant e^{\beta Y_{\tau^{k}_{\delta}}}+\int_{\delta}^{\tau^{k}_{\delta}}|\beta|e^{\beta y^{n}_{s}}|g^{n}_{s}|ds-\int_{\delta}^{\tau^{k}_{\delta}}\beta e^{\beta y^{n}_{s}}(Z^{n}_{s}-z)\cdot dB_{s}$ (5.17)
$\leqslant e^{\beta Y_{\tau^{k}_{\delta}}}+(\mu+\nu)(1+2|z|^{2})|\beta|\int_{\delta}^{\tau^{k}_{\delta}}e^{\beta y^{n}_{s}}ds+2(\mu+\nu)|\beta|\int_{\delta}^{\tau^{k}_{\delta}}e^{\beta y^{n}_{s}}|Z^{n}_{s}-z|^{2}ds$
$\ \ \ \ \ \ -\int_{\delta}^{\tau^{k}_{\delta}}\beta e^{\beta y^{n}_{s}}(Z^{n}_{s}-z)\cdot dB_{s},$
which together with (5.9) implies
$\left(\frac{1}{2}\beta^{2}-2(\mu+\nu)|\beta|\right)$ $E\left[\int_{\delta}^{\tau^{k}_{\delta}}e^{\beta y^{n}_{s}}|Z^{n}_{s}-z|^{2}ds|{\cal{F}}_{\delta}\right]\leqslant e^{|\beta|C}(1+(\mu+\nu)(1+2|z|^{2})|\beta|T).$ (5.18)
Choose $\beta<-4(\mu+\nu)$ such that $\frac{1}{2}\beta^{2}-2(\mu+\nu)|\beta|=1$. By (5.9) and Fatou’s lemma, when $k$ tends to $\infty$, we have
$E\left[\int_{\delta}^{T}|Z^{n}_{s}-z|^{2}ds|{\cal{F}}_{\delta}\right]\leqslant e^{2|\beta|C}(1+(\mu+\nu)(1+2|z|^{2})|\beta|T).$ (5.19)
From this and the energy inequality for BMO martingales (see [22, Page 29]), we conclude that for all $m\geqslant 1$,
$\sup_{n\geqslant 1}E\left[\left(\int_{0}^{T}|Z_{s}^{n}|^{2}ds\right)^{m}\right]\leqslant\sup_{n\geqslant 1}m!\|Z^{n}_{s}\|^{2m}_{BMO}<\infty.$ (5.20)
By (5.12), (5.9) and (5.13), we have
$|A_{T}^{n}|^{2}$ $\leqslant\left(2C+\int_{0}^{T}|g_{s}^{n}|ds+\left|\int_{0}^{T}(Z^{n}_{s}-z)\cdot dB_{s}\right|\right)^{2}$
$\leqslant 3\left(4C^{2}+2(\mu+\nu)^{2}T^{2}+2(\mu+\nu)^{2}\left(\int_{0}^{T}|Z_{s}^{n}|^{2}ds\right)^{2}+\left|\int_{0}^{T}(Z^{n}_{s}-z)\cdot dB_{s}\right|^{2}\right).$
Taking expectation and then by BDG inequality and (5.20), we have
$\sup_{n\geqslant 1}E|A_{T}^{n}|^{2}<\infty.$ (5.21)
Then, by (5.9), dominated convergence theorem, (5.10) and (5.21), we have
$E\int_{0}^{T}\lim_{n\rightarrow\infty}|Y_{s}-y_{s}^{n}|ds=\lim_{n\rightarrow\infty}E\int_{0}^{T}|Y_{s}-y_{s}^{n}|ds\leqslant\lim_{n\rightarrow\infty}\frac{1}{n}\sup_{n\geqslant 1}E\left[|A_{T}^{n}|\right]=0,$
which implies that
$y^{n}_{t}\nearrow Y_{t},\ dt\times dP\text{-}a.e.,\quad\text{as}\quad n\rightarrow\infty.$ (5.22)
By Proposition 5.3(ii) and (A), we have for all $0\leqslant s\leqslant t\leqslant T,$
$-y_{s}^{n}-z\cdot B_{s}\leqslant-\rho_{s}(-y_{t}^{n}-z\cdot B_{t})\leqslant-\mathcal{E}^{-\mu,-\nu}[y_{t}^{n}+z\cdot B_{t}|\mathcal{F}_{s}]=\mathcal{E}^{\mu,\nu}[-y_{t}^{n}-z\cdot B_{t}|\mathcal{F}_{s}],$
which implies that $y_{t}^{n}-z\cdot B_{t}$ is a $g^{\mu,\nu}$-submartingale. It follows from Lemma 2.7 that there exists a pair $(\tilde{Z}^{n}_{s},\tilde{A}_{s})\in H^{2}_{d}\times\mathcal{A},$ such that
$y_{t}^{n}+z\cdot B_{t}=Y_{T}+z\cdot B_{T}+\tilde{A}_{T}^{n}-\tilde{A}_{t}^{n}+\int_{t}^{T}g^{-\mu,-\nu}(-\tilde{Z}_{s}^{n})ds-\int_{t}^{T}-\tilde{Z}_{s}^{n}\cdot dB_{s},\quad t\in[0,T],$ (5.23)
Thus, $\tilde{A}_{t}^{n}$ is continuous. Comparing (5.23) with (5.12), we have $dt\times dP$-$a.e.$, $\tilde{Z}_{s}^{n}=Z_{s}^{n}$. Applying Itô’s formula to $\exp(-2\nu(y_{t}^{n}+z\cdot B_{t}))$, we have
$\tilde{y}_{t}^{n}=-\exp(-2\nu(Y_{T}+z\cdot B_{T}))+\tilde{a}_{T}^{n}-\tilde{a}_{t}^{n}+\int_{t}^{T}-\mu|\tilde{z}_{s}^{n}|ds-\int_{t}^{T}\tilde{z}_{s}^{n}\cdot dB_{s},\quad t\in[0,T],$ (5.24)
with $\tilde{y}_{t}^{n}=-\exp(-2\nu(y_{t}^{n}+z\cdot B_{t})),\ \tilde{z}_{t}^{n}=2\nu\exp(-2\nu(y_{t}^{n}+z\cdot B_{t}))Z_{s}^{n}$ and $\tilde{a}_{t}^{n}=2\nu\int_{0}^{t}\exp(-2\nu(y_{s}^{n}+z\cdot B_{s}))d\tilde{A}^{n}_{s}$. Then by (5.9), we have
$\sup_{n\geqslant 1}E\left[\sup_{t\in[0,T]}|\tilde{y}_{t}^{n}|^{2}\right]$ $=\sup_{n\geqslant 1}E\left[\sup_{t\in[0,T]}\exp(-4\nu(y_{t}^{n}+z\cdot B_{t}))\right]$
$\leqslant\exp(4\nu C)E\left[\sup_{t\in[0,T]}\exp(-4\nu z\cdot B_{t}))\right]<\infty.$ (5.25)
From this and Bouchard et al. [6, Theorem 2.1], we further have
$\sup_{n\geqslant 1}E\left(\int_{0}^{T}|\tilde{z}_{s}^{n}|^{2}ds+|\tilde{a}_{T}^{n}|^{2}\right)<\infty.$ (5.26)
By (5.9), there exists a bounded progressively measurable process $\tilde{Y}_{t}$ such that for all $t\in[0,T]$, $y^{n}_{t}\nearrow\tilde{Y}_{t}$, as $n\rightarrow\infty$. This implies that for all $t\in[0,T]$, $\tilde{y}_{t}^{n}\nearrow-\exp(-2\nu(\tilde{Y}_{t}+z\cdot B_{t}))$, as $n\rightarrow\infty$. Then, in view of (5.24)-(5.26), by [29, Theorem 2.4], we deduce that $\exp(-2\nu(\tilde{Y}_{t}+z\cdot B_{t}))$ is RCLL and thus $\tilde{Y}_{t}$ is RCLL. This with (5.22) implies that for all $t\in[0,T]$, $\tilde{Y}_{t}=Y_{t}$. Then by Dini’s theorem, we have
$\lim_{n\rightarrow\infty}\sup_{t\in[0,T]}|y^{n}_{t}-Y_{t}|\rightarrow 0.$
In view of this and (5.9), by dominated convergence theorem, we obtain (5.15).
Applying Itô’s formula to $|y^{n}_{t}-y^{m}_{t}|^{2}$ and (5.13), we have
$E$ $|y^{n}_{0}-y^{m}_{0}|^{2}+E\int_{0}^{T}|Z^{n}_{s}-Z^{m}_{s}|^{2}ds$
$=2E\int_{0}^{T}(y^{n}_{s}-y^{m}_{s})(g^{n}_{s}-g^{m}_{s})ds+2E\int_{0}^{T}(y^{n}_{s}-y^{m}_{s})d(A^{n}_{s}-A^{m}_{s})$
$\leqslant 2(\mu+\nu+1)E\left[\sup_{t\in[0,T]}|y^{n}_{t}-y^{m}_{t}|\left(\int_{0}^{T}(2+|Z^{n}_{s}|^{2}+|Z^{m}_{s}|^{2})ds+(|A^{n}_{T}|+|A^{m}_{T}|)\right)\right]$
$\leqslant 2(\mu+\nu+1)\sqrt{E\left[\sup_{t\in[0,T]}|y^{n}_{t}-y^{m}_{t}|^{2}\right]}\sqrt{2E\left[\left(\int_{0}^{T}(2+|Z^{n}_{s}|^{2}+|Z^{m}_{s}|^{2})ds\right)^{2}\right]+2E|A^{n}_{T}+A^{m}_{T}|^{2}},$
which with (5.15), (5.20) and (5.21) implies that $\{Z^{n}_{t}\}_{n\geqslant 1}$ is a Cauchy sequence in ${\cal{H}}^{2}_{d}.$
By (5.14) and (5.13), we have
$g_{t}^{n}-g_{t}^{m}$ $=g_{t}^{n}-\theta g_{t}^{m}-(1-\theta)g_{t}^{m}$
$\leqslant(\mu+\nu)(1-\theta)\left(1+\left|Z^{n}_{t}+\frac{\theta}{1-\theta}(Z^{n}_{t}-Z^{m}_{t})\right|^{2}\right)+(1-\theta)|g_{t}^{m}|$
$\leqslant(\mu+\nu)(1-\theta)\left(1+2|Z^{n}_{t}|^{2}+2\left(\frac{\theta}{1-\theta}\right)^{2}|Z^{n}_{t}-Z^{m}_{t}|^{2}\right)+\mu(1-\theta)(1+|Z^{m}_{t}|^{2})$
$\leqslant 2(\mu+\nu)(1-\theta)\left(1+|Z^{n}_{t}|^{2}+|Z^{m}_{t}|^{2}\right)+\frac{2(\mu+\nu)\theta^{2}}{1-\theta}|Z^{n}_{t}-Z^{m}_{t}|^{2}.$
By interchanging $g_{t}^{n}$ and $g_{t}^{m}$, we deduce that
$E\int_{0}^{T}|g_{t}^{n}-g_{t}^{m}|dt\leqslant 4(\mu+\nu)(1-\theta)\sup_{n\geqslant 1}E\int_{0}^{T}\left(1+|Z^{n}_{t}|^{2}\right)dt+\frac{2(\mu+\nu)\theta^{2}}{1-\theta}E\int_{0}^{T}|Z^{n}_{t}-Z^{m}_{t}|^{2}dt.$
Since (5.20) holds and $\{Z^{n}_{t}\}_{n\geqslant 1}$ is a Cauchy sequence in ${\cal{H}}^{2}_{d}$, letting $n,m\rightarrow\infty$, and then $\theta\rightarrow 1$, we get that $\{g^{n}_{t}\}_{n\geqslant 1}$ is a Cauchy sequence in ${\cal{H}}^{1}_{1}.$ We obtain (5.16). ∎
By (5.12) and Proposition 5.5, we deduce that for all $t\in[0,T]$, $\{A^{n}_{t}\}_{n\geqslant 1}$ is a Cauchy sequence in $L^{1}({\cal{F}}_{t}).$ We denote the limit of $\{A^{n}_{t}\}_{n\geqslant 1}$ in $L^{1}({\cal{F}}_{t})$ by $A_{t}$. Then we deduce that $A_{t}$ is an increasing and continuous process with $A_{0}=0$.
Since $\{Z^{n}_{t}\}_{n\geqslant 1}$ is a Cauchy sequence in ${\cal{H}}^{2}_{d}$ and $\{\int_{0}^{T}Z^{n}_{t}\cdot dB_{t}\}_{n\geqslant 1}$ is a Cauchy sequence in $L^{2}({\cal{F}}_{T})$, by Kobylanski [24, Lemma 2.5], we can find a subsequence of $\{Z^{n}_{t}\}_{n\geqslant 1}$ still denoted by $\{Z^{n}_{t}\}_{n\geqslant 1}$ such that $\sup_{n\geqslant 1}|Z_{t}^{n}|\in{\cal{H}}^{2}_{d}$ and $\sup_{n\geqslant 1}\left|\int_{0}^{T}Z^{n}_{t}\cdot dB_{t}\right|\in L^{2}({\cal{F}}_{T})$. From this, (5.12), (5.9) and (5.13), we have
$E[\sup_{n\geqslant 1}A_{T}^{n}]$ $\leqslant 2C+E\left[\int_{0}^{T}\sup_{n\geqslant 1}|g_{s}^{n}|ds\right]+E\left[\sup_{n\geqslant 1}\left|\int_{0}^{T}Z^{n}_{s}\cdot dB_{s}\right|\right]$
$\leqslant 2C+(\mu+\nu)E\left[\int_{0}^{T}\sup_{n\geqslant 1}(1+|Z^{n}_{s}|^{2})ds\right]+E\left[\sup_{n\geqslant 1}\left|\int_{0}^{T}Z^{n}_{s}\cdot dB_{s}\right|\right]<\infty.$
Set
$\sigma_{k}:=\inf\left\{t\geqslant 0,E\left[\sup_{n\geqslant 1}A_{T}^{n}|{\cal{F}}_{t}\right]\geqslant k\right\}\wedge T,\ \ k>0.$
By (5.11) and Proposition 5.2, we get that for all $t\in[0,T]$ and $k>0,$
$\rho_{t}(-y^{n}_{\sigma_{k}}-z\cdot B_{\sigma_{k}}-A^{n}_{\sigma_{k}})=y^{n}_{{\sigma_{k}}\wedge t}+z\cdot B_{{\sigma_{k}}\wedge t}+A^{n}_{{\sigma_{k}}\wedge t}.$
Since $A^{n}_{t}$ is increasing, we have $\sup_{n\geqslant 1}A^{n}_{\sigma_{k}}\leqslant E\left[\sup_{n\geqslant 1}A_{T}^{n}|{\cal{F}}_{\sigma_{k}}\right]\leqslant k$. Hence, by (5.9), (5.15) and Proposition 4.2(i), we get that for all $t\in[0,T]$ and $k>0,$
$\rho_{t}(-Y_{\sigma_{k}}-z\cdot B_{\sigma_{k}}-A_{\sigma_{k}})=Y_{{\sigma_{k}}\wedge t}+z\cdot B_{{\sigma_{k}}\wedge t}+A_{{\sigma_{k}}\wedge t}.$
Moreover, if $A_{T}$ is bounded, then letting $k\rightarrow\infty$, by Proposition 4.2(i) again, we get that for all $t\in[0,T],$
$\rho_{t}(-Y_{T}-z\cdot B_{T}-A_{T})=Y_{t}+z\cdot B_{t}+A_{t}.$
The proof is complete. ∎
6 The proofs of Theorem 3.3 and Corollary 3.4
The proof of Theorem 3.3:.
Step 1. For $z\in\textbf{R}^{d}$, we have
$g^{\mu,\nu}(z)t+z\cdot B_{t}=-g^{\mu,\nu}(z)T+z\cdot B_{T}+\int_{t}^{T}g^{\mu,\nu}(z)ds-\int_{t}^{T}z\cdot dB_{s},\quad t\in[0,T],$ (6.1)
which implies that $g^{\mu,\nu}(z)t+z\cdot B_{t}$ is a $g^{\mu,\nu}$-martingale. From (A), it follows that $g^{\mu,\nu}(z)t+z\cdot B_{t}$ is a $\rho$-supermartingale. Then by Theorem 5.4, there exists a process $A_{t}^{z}\in{\cal{A}}$ and an increasing series $\{\tau_{n}\}_{n\geqslant 1}\subset\mathcal{T}_{0,T}$ satisfying $\tau_{n}\nearrow T$ as $n\rightarrow\infty$, such that for all $n\geqslant 1$ and $t\in[0,T],$
$\rho_{t}(g^{\mu,\nu}(z)\tau_{n}-z\cdot B_{\tau_{n}}-A_{\tau_{n}}^{z})=-g^{\mu,\nu}(z)({t\wedge\tau_{n}})+zB_{t\wedge\tau_{n}}+A_{t\wedge\tau_{n}}^{z}.$
From Proposition 4.3, it follows that for all $n\geqslant 1$, there exists a pair $(g_{n}(s,z),Z^{z,n}_{s})\in H_{1}^{1}\times H_{d}^{2}$ such that
$g^{\mu,\nu}(z)({t\wedge\tau_{n}})+z\cdot B_{t\wedge\tau_{n}}+A_{t\wedge\tau_{n}}^{z}=-g^{\mu,\nu}(z){\tau_{n}}+z\cdot B_{\tau_{n}}+A_{\tau_{n}}^{z}+\int_{t\wedge\tau_{n}}^{\tau_{n}}g_{n}(s,z)ds-\int_{t\wedge\tau_{n}}^{\tau_{n}}Z^{z,n}_{s}\cdot dB_{s}.$ (6.2)
Comparing (6.1) and (6.2), we have for all $n\geqslant 1$, $dt\times dP$-$a.e.,$
$Z^{z,n}_{t}=z\quad\text{and}\quad|g_{n}(t,z)|\leqslant\mu|z|+\nu|z|^{2},\quad t\in[0,\tau_{n}],$ (6.3)
and thus,
$A_{\tau_{n}}^{z}=\int_{0}^{\tau_{n}}g^{\mu,\nu}(z)ds-\int_{0}^{\tau_{n}}g_{n}(s,z)ds\leqslant 2T(\mu|z|+\nu|z|^{2})<\infty,$
which implies that $A_{T}^{z}$ is bounded. Then by Theorem 5.4 again, we get that for all $t\in[0,T],$
$\rho_{t}(g^{\mu,\nu}(z)T-z\cdot B_{T}-A_{T}^{z})=-g^{\mu,\nu}(z)t+z\cdot B_{t}+A_{t}^{z}.$ (6.4)
By Proposition 4.3 again and a similar argument as in (6.3), we deduce that there exists a process $g(t,z)\in H_{1}^{1}$ such that
$g^{\mu,\nu}(z)t+z\cdot B_{t}+A_{t}^{z}=-g^{\mu,\nu}(z)T+z\cdot B_{T}+A_{T}^{z}+\int_{t}^{T}g(s,z)ds-\int_{t}^{T}z\cdot dB_{s},$ (6.5)
where $dt\times dP$-$a.e.,$ $|g(t,z)|\leqslant\mu|z|+\nu|z|^{2}$ and for all $\tilde{z}\in\textbf{R}^{d}$ and $\theta\in(0,1)$,
$g(t,z)-\theta g(t,\tilde{z})\leqslant\mu|z-\tilde{z}|+\frac{\nu}{1-\theta}|z-\tilde{z}|^{2}.$
This implies that $g\in{\cal{G}}^{\mu,\nu}\cap{\cal{G}}^{\mu,\nu}_{\theta}$. Moreover, by (6.5), Proposition 4.1(i), (6.4) and Proposition 5.2, we get that for all $z\in\mathbf{R}^{d}$, $\tau\in\mathcal{T}_{0,T}$ and $0\leqslant r\leqslant t\leqslant T$,
$\rho_{\tau\wedge r}\left(\int_{\tau\wedge r}^{\tau\wedge t}g(s,z)ds-\int_{\tau\wedge r}^{\tau\wedge t}z\cdot dB_{s}\right)$
$=\rho_{\tau\wedge r}(g^{\mu,\nu}(z)({\tau\wedge t})-z\cdot B_{\tau\wedge t}^{z}-A_{\tau\wedge t}^{z}-(g^{\mu,\nu}(z)({\tau\wedge r})-z\cdot B_{\tau\wedge r}^{z}-A_{\tau\wedge r}^{z}))$
$=\rho_{\tau\wedge r}(g^{\mu,\nu}(z)({\tau\wedge t})-z\cdot B_{\tau\wedge t}^{z}-A_{\tau\wedge t}^{z})+(g^{\mu,\nu}(z)({\tau\wedge r})-z\cdot B_{\tau\wedge r}^{z}-A_{\tau\wedge r}^{z})$
$=0.$
Step 2. In view of the consequence of Step 1 and Proposition 4.2(ii), by similar arguments as in [36, (4.8)], for any $\tau\in\mathcal{T}_{0,T}$, $r\in[0,T]$ and ${\cal{F}}_{\tau\wedge r}$-measurable simple random variable $\eta$ (i.e., $\eta=\sum^{m}_{j=1}z_{j}1_{A_{j}}$ where for all $1\leqslant j\leqslant m$, $z_{j}\in{\bf{R}}^{d}$ and $\{A_{j}\}_{j=1}^{m}$ is a partition of $\Omega$ such that for all $1\leqslant j\leqslant m$, $A_{j}\in{\cal{F}}_{\tau\wedge r}$), we have
$\rho_{\tau\wedge r}\left(\int_{\tau\wedge r}^{\tau\wedge t}g(s,\eta)ds-\int_{\tau\wedge r}^{\tau\wedge t}\eta\cdot dB_{s}\right)=0,\quad\forall t\geqslant r.$ (6.6)
Let $\cal{O}$} be the set of all the simple step processes defined as
$\eta_{t}=\sum^{l}_{i=0}\sum^{m_{i}}_{j=1}z_{ij}1_{[t_{i},t_{i+1})\times A_{ij}}(t,\omega),\ \ (t,\omega)\in[0,T]\times\Omega,$
where $0=t_{0}<t_{1}<\cdots<t_{l}<t_{l+1}=T$, $z_{ij}\in{\bf{R}}^{d}$ and $\{A_{ij}\}_{j=1}^{m_{i}}$ is a partition of $\Omega$ such that for all $0\leqslant i\leqslant l$ and $0\leqslant j\leqslant m_{i}$, $A_{ij}\in{\cal{F}}_{t_{i}}$. For any $\eta_{s}\in{\cal{O}}$, $\tau\in\mathcal{T}_{0,T}$ and $0\leqslant r\leqslant T$ with $r\in[t_{i},t_{i+1})$ for some $0\leqslant i\leqslant l$, by setting $r_{i}=r$ and $r_{k}=t_{k},$ $i<k\leqslant l+1$, we have
$\rho_{\tau\wedge r}\left(\int_{\tau\wedge r}^{\tau}g(s,\eta_{s})ds-\int_{\tau\wedge r}^{\tau}\eta_{s}\cdot dB_{s}\right)$
$=\rho_{\tau\wedge r}\left(\sum^{l}_{k=i}\left(\int_{\tau\wedge r_{k}}^{\tau\wedge r_{k+1}}g(s,\eta_{s})ds-\int_{\tau\wedge r_{k}}^{\tau\wedge r_{k+1}}\eta_{s}\cdot dB_{s}\right)\right)$
$=\rho_{\tau\wedge r}\left(-\rho_{\tau\wedge r_{l}}\left(\int_{\tau\wedge r_{l}}^{\tau}g(s,\eta_{\tau\wedge r_{l}})ds-\int_{\tau\wedge r_{l}}^{\tau}\eta_{\tau\wedge r_{l}}\cdot dB_{s}\right)\right$.
[…]
$=\rho_{\tau\wedge r}\left(\sum^{l-1}_{k=i}\left(\int_{\tau\wedge r_{k}}^{\tau\wedge r_{k+1}}g(s,\eta_{s})ds-\int_{\tau\wedge r_{k}}^{\tau\wedge r_{k+1}}\eta_{s}\cdot dB_{s}\right)\right)\quad\quad\text{by}\ (\ref{6.6})$
$\cdots$
$=\rho_{\tau\wedge r}\left(\int_{\tau\wedge r}^{\tau\wedge r_{i+1}}g(s,\eta_{\tau\wedge r})ds-\int_{\tau\wedge r}^{\tau\wedge r_{i+1}}\eta_{\tau\wedge r}\cdot dB_{s}\right)$
$=0.\quad\quad\text{by}\ (\ref{6.6})$ (6.7)
Since $g(t,z)\in{\cal{G}}^{\mu,\nu}\cap{\cal{G}}^{\mu,\nu}_{\theta}$, by [12, Theorem 2.6(i)], for $\xi\in\mathcal{L}^{\exp}_{T},$ the following BSDE:
$Y_{t}=-\xi+\int_{t}^{T}g(s,Z_{s})ds-\int_{t}^{T}Z_{s}\cdot dB_{s},\quad t\in[0,T].$ (6.8)
has a unique solution $(Y_{t},Z_{t})\in\mathcal{L}^{\exp}_{\mathcal{F}}\times\bigcap_{r>1}\mathcal{H}_{d}^{r}.$ It is clear that there exists a sequence $\{Z^{n}_{t}\}_{n\geqslant 1}\subset{\cal{O}}$ such that $Z^{n}_{t}\rightarrow Z_{t}$ in ${\cal{H}}^{2}_{d}$ as $n\rightarrow\infty.$ Since
$\left|\sup_{t\in[0,T]}\left|\int_{0}^{t}Z^{n}_{s}\cdot dB_{s}\right|-\sup_{t\in[0,T]}\left|\int_{0}^{t}Z_{s}\cdot dB_{s}\right|\right|\leqslant\sup_{t\in[0,T]}\left|\int_{0}^{t}(Z^{n}_{s}-Z_{s})\cdot dB_{s}\right|,$
by BDG inequality, we have $\sup_{t\in[0,T]}\left|\int_{0}^{t}Z^{n}_{s}\cdot dB_{s}\right|\rightarrow\sup_{t\in[0,T]}\left|\int_{0}^{t}Z_{s}\cdot dB_{s}\right|$ in $L^{2}({\cal{F}}_{T})$ as $n\rightarrow\infty.$ Then by [24, Lemma 2.5], we can find a subsequence of $\{Z^{n}_{t}\}_{n\geqslant 1}$ still denoted by $\{Z^{n}_{t}\}_{n\geqslant 1}$ such that $\sup_{n\geqslant 1}|Z^{n}_{t}|\in{\cal{H}}^{2}_{d}$ and $\sup_{n\geqslant 1}\sup_{t\in[0,T]}\left|\int_{0}^{t}Z^{n}_{s}\cdot dB_{s}\right|\in L^{2}({\cal{F}}_{T}).$
For $k>0$, we define
$\sigma_{k}:=\inf\left\{t\geqslant 0:\int_{0}^{t}\sup_{n\geqslant 1}|Z^{n}_{s}|^{2}ds+\sup_{n\geqslant 1}\sup_{r\in[0,t]}\left|\int_{0}^{r}Z^{n}_{s}\cdot dB_{s}\right|\geqslant k\right\}\wedge T.$
By Début theorem, we get that for any $k>0$, $\sigma_{k}$ is a stopping time. Since $g(t,z)\in{\cal{G}}^{\mu,\nu}\cap{\cal{G}}^{\mu,\nu}_{\theta}$, by Remark 2.1 and dominated convergence theorem, we deduce that for all $k>0$, there exists a subsequence of $\{Z^{n}_{t}\}_{n\geqslant 1}$ still denoted by $\{Z^{n}_{t}\}_{n\geqslant 1}$ such that
$\int^{\sigma_{k}}_{\sigma_{k}\wedge t}g(s,Z^{n}_{s})ds-\int^{\sigma_{k}}_{\sigma_{k}\wedge t}Z^{n}_{s}\cdot dB_{s}\rightarrow\int^{\sigma_{k}}_{\sigma_{k}\wedge t}g(s,Z_{s})ds-\int^{\sigma_{k}}_{\sigma_{k}\wedge t}Z_{s}\cdot dB_{s},\quad dP\text{-}a.s.,$
as $\rightarrow\infty.$ Moreover, since $\sup_{r\in[0,t]}\left|\int_{0}^{r}Z^{n}_{s}\cdot dB_{s}\right|$ is continuous in $t$ and $\sup_{n\geqslant 1}\sup_{r\in[0,t]}\left|\int_{0}^{r}Z^{n}_{s}\cdot dB_{s}\right|$ is increasing in $t$, we can deduce that $\sup_{n\geqslant 1}\sup_{r\in[0,t]}\left|\int_{0}^{r}Z^{n}_{s}\cdot dB_{s}\right|$ is left-continuous in $t$. Thus, for all $k>0$, we have
$\sup_{n\geqslant 1}\left|\int^{\sigma_{k}}_{\sigma_{k}\wedge t}g(s,Z^{n}_{s})ds+\int^{\sigma_{k}}_{\sigma_{k}\wedge t}Z^{n}_{s}\cdot dB_{s}\right|$
$\leqslant 2\int^{\sigma_{k}}_{0}(\mu+\nu)(1+\sup_{n\geqslant 1}|Z^{n}_{s}|^{2})ds+2\sup_{n\geqslant 1}\sup_{r\in[0,\sigma_{k}]}\left|\int^{r}_{0}Z^{n}_{s}\cdot dB_{s}\right|$
$\leqslant 2(\mu+\nu)(T+k)+2k.$
By the two equations above, (6) and Proposition 4.2(i), we get that for all $k>0$,
$\rho_{\sigma_{k}\wedge t}\left(\int^{\sigma_{k}}_{\sigma_{k}\wedge t}g(s,Z_{s})ds-\int^{\sigma_{k}}_{\sigma_{k}\wedge t}Z_{s}\cdot dB_{s}\right)=0.$
Then by (6.8), Proposition 4.1(i) and Definition 2.2(i), we get that for all $k>0$ and $t\in[0,T]$,
$\rho_{\sigma_{k}\wedge t}(-Y_{\sigma_{k}})=Y_{\sigma_{k}\wedge t}=\mathcal{E}^{g}[-\xi|{\cal{F}}_{\sigma_{k}\wedge t}].$
From this and Proposition 4.2(i), we deduce that for all $\xi\in\mathcal{L}^{\exp}_{T},$ (3.2) holds. Moreover, by Lemma 2.3(vii)(viii), we further get that $g$ is a unique function in ${\cal{G}}^{\mu,\nu}\cap{\cal{G}}_{conv}$ such that for all $\xi\in\mathcal{L}^{\exp}_{T},$ (3.2) holds.
Step 3. In view of (3.2), to prove (3.3), we only need to show that for all $\xi\in\mathcal{L}^{\exp}_{T},$
$\mathcal{E}^{g}[-\xi|{\cal{F}}_{t}]=\operatorname{ess\,sup}_{q\in\mathcal{Q}_{\xi,f}}E_{Q^{q}}\left[-\xi-\int_{t}^{T}f(s,q_{s})ds|\mathcal{F}_{t}\right],\quad\forall t\in[0,T],$ (6.9)
where $\mathcal{Q}_{\xi,f}:=\{q_{t}\in H_{d}^{2}:\theta^{q}_{t}:=\exp\{\int_{0}^{t}q_{s}\cdot dB_{s}-\frac{1}{2}\int_{0}^{t}|q_{s}|^{2}ds\}$ is a uniformly integrable martingale such that $E_{Q^{q}}[|\xi|+\int_{0}^{T}|f(s,q_{s})|ds]<\infty$ with $dQ^{q}=\theta^{q}_{T}dP\}.$ (6.9) can be derived from the proof of [13, Theorem 3.1(i)]. For convenience, we show its proof in the Appendix, where we can also find that the supremum in (3.3) is attained by any $q_{t}\in H^{2}_{d}$ such that $dt\times dP$-$a.e.,$ $q_{t}\in\partial g(t,Z_{t}^{-\xi,g})$. The proof is complete. ∎
The proof of Corollary 3.4:.
Step 1. By Theorem 3.3, there exists a unique $g\in{\cal{G}}^{\mu,\nu}\cap{\cal{G}}_{conv}$ such that for all $\xi\in\mathcal{L}^{\exp}_{T}$,
$\rho_{t}(\xi)={\cal{E}}^{g}[-\xi|{\cal{F}}_{t}],\quad t\in[0,T].$ (6.10)
Since $\rho$ satisfies (r5) and (r6) on $\mathcal{L}^{\exp}_{T}$, by Lemma 2.3(ix), we get that $g\in{\cal{G}}^{\mu,\nu}\cap{\cal{G}}_{subl}$. This implies that for all $z\in\mathbf{R}^{d}$ and $\beta>0$, $\beta|g(z)|=|g(\beta z)|\leqslant\beta\mu|z|+\beta^{2}\nu|z|^{2}$, and thus
$|g(z)|\leqslant\mu|z|+\beta\nu|z|^{2}.$
Letting $\beta\rightarrow 0$, we have $g\in{\cal{G}}^{\mu,0}\cap{\cal{G}}_{subl}$.
Since $g\in{\cal{G}}^{\mu,0}$, we have, $dt\times dP$-$a.e.$,
$f(t,x):=\sup_{z\in\mathbf{R}^{d}}\{x\cdot z-g(t,z)\}\geqslant\sup_{z\in\mathbf{R}^{d}}\{x\cdot z-\mu|z|\}=\infty 1_{\{|x|>\mu\}},\quad\forall x\in\mathbf{R}^{d}.$ (6.11)
For $x\in\mathbf{R}^{d}$, if for all $z\in\mathbf{R}^{d}$, $x\cdot z\leqslant g(t,z)$, which is equivalent to $x\in\partial g(t,0)$ (since $g(t,0)\equiv 0$), then we have $dt\times dP$-$a.e.$,
$f(t,x)=\sup_{z\in\mathbf{R}^{d}}\{x\cdot z-g(t,z)\}=0.$ (6.12)
and if for some $z\in\mathbf{R}^{d}$, $x\cdot z>g(t,z)$, then since $g\in{\cal{G}}_{subl}$, we have, $dt\times dP$-$a.e.$, for all $\beta>0$,
$f(t,x)\geqslant\beta x\cdot z-g(t,\beta z)=\beta(x\cdot z-g(t,z))\rightarrow\infty,\quad\text{as}\ \beta\rightarrow\infty.$ (6.13)
By (6.11)-(6.13), we get that $dt\times dP$-$a.e.$,
$f(t,x)=\infty 1_{\{x\notin\partial g(t,0)\}},\quad\forall x\in\mathbf{R}^{d}.$ (6.14)
and for all $x\in\partial g(t,0)$, $|x|\leqslant\mu.$
By [12, Theorem 2.6(i)], for $\xi\in\mathcal{L}^{\exp}_{T},$ the following BSDE:
$Y_{t}=-\xi+\int_{t}^{T}g(s,Z_{s})ds-\int_{t}^{T}Z_{s}\cdot dB_{s},\quad t\in[0,T].$ (6.15)
has a unique solution $(Y_{t},Z_{t})\in\mathcal{L}^{\exp}_{\mathcal{F}}\times\bigcap_{r>1}\mathcal{H}_{d}^{r}.$ By measurable selection theorem (or [3, Lemma 7.5]), we can find a progressively measurable process $\tilde{q}_{t}\in\partial g(t,Z_{t})$, $dt\times dP$-$a.e.$ This implies that $dt\times dP$-$a.e.,$ $f(t,\tilde{q}_{t})=\tilde{q}_{t}\cdot Z_{t}-g(t,Z_{t}),$ and then by (6.14), we have $dt\times dP$-$a.e.,$ $\tilde{q}_{t}\in\partial g(t,0)$, $|\tilde{q}_{t}|\leqslant\mu$ and $f(t,\tilde{q}_{t})=0$. Set $dQ^{\tilde{q}}/dP:=\exp(\int_{0}^{T}\tilde{q}_{s}\cdot dB_{s}-\frac{1}{2}\int_{0}^{T}|\tilde{q}_{s}|^{2}ds)$. Clearly, $Q^{\tilde{q}}\in\mathcal{Q}_{g}$. Since $dt\times dP$-$a.e.,$ $|\tilde{q}_{t}|\leqslant\mu$, we can deduce that for all $p\geqslant 2$, $dQ^{\tilde{q}}/dP\in L^{p}(\mathcal{F}_{T})$. Thus, by Hölder’s inequality, we have for all $t\in[0,T]$, $E_{Q^{\tilde{q}}}[|Y_{t}|]<\infty$ and then by (6.15), we have
$Y_{t}$ $=E_{Q^{\tilde{q}}}\left[-\xi+\int_{t}^{T}g(s,Z_{s})ds-\int_{t}^{T}Z_{s}\cdot dB_{s}|\mathcal{F}_{t}\right]$
$=E_{Q^{\tilde{q}}}\left[-\xi-\int_{t}^{T}Z_{s}\cdot dB^{\tilde{q}}_{s}|\mathcal{F}_{t}\right]$
$=E_{Q^{\tilde{q}}}[-\xi|\mathcal{F}_{t}],\quad\forall t\in[0,T],$ (6.16)
where $B^{\tilde{q}}_{t}:=B_{t}-\int_{0}^{t}{\tilde{q}}_{s}ds$ is a standard Brownian motion under the probability $Q^{\tilde{q}}$.
For any $q_{t}\in\mathcal{Q}_{g},$ by (6.14) and (6.11), we have $dt\times dP$-$a.e.,$ $0=f(t,q_{t})\geqslant q_{t}\cdot Z_{t}-g(t,Z_{t}).$ Then by a similar argument as in (6), we have
$Y_{t}\geqslant E_{Q^{q}}[-\xi|\mathcal{F}_{t}],\quad t\in[0,T].$
From this, (6) and Definition 2.2(i), it follows that for all $\xi\in\mathcal{L}^{\exp}_{T}$,
${\cal{E}}^{g}[-\xi|{\cal{F}}_{t}]=\sup_{Q\in\mathcal{Q}_{g}}E_{Q}[-\xi|\mathcal{F}_{t}],\quad t\in[0,T].$ (6.17)
This with (6.10) gives (3.4).
If there exists another $\tilde{g}\in{\cal{G}}^{\mu,0}\cap{\cal{G}}_{subl}$ such that (3.4) holds, then by the same arguments as in (6.11)-(6.17), we deduce that for all $\xi\in\mathcal{L}^{\exp}_{T}$ and $t\in[0,T]$,
${\cal{E}}^{\tilde{g}}[-\xi|{\cal{F}}_{t}]=\sup_{Q\in\mathcal{Q}_{\tilde{g}}}E_{Q}[-\xi|\mathcal{F}_{t}]=\rho_{t}(\xi)=\sup_{Q\in\mathcal{Q}_{g}}E_{Q}[-\xi|\mathcal{F}_{t}]={\cal{E}}^{g}[-\xi|{\cal{F}}_{t}].$
This with Lemma 2.3(vi) implies that $dt\times dP$-$a.e.,$ for all $z\in\mathbf{R}^{d}$, $g(t,z)=\tilde{g}(t,z).$
Step 2. Let us further assume that for all $\xi\in L^{2}(\mathcal{F}_{T})$, $\rho$ satisfies (3.1) with $\nu=0$. Then, by Proposition 4.1(ii) and (A), for all $\xi,\eta\in L^{2}(\mathcal{F}_{T})$ and $\theta\in(0,1)$, we have
$\rho_{t}(\xi)-\theta\rho_{t}(\eta)$ $\leqslant(1-\theta)\mathcal{E}^{\mu,0}\left[\frac{-\xi+\theta\eta}{1-\theta}|\mathcal{F}_{t}\right]$
$=\mathcal{E}^{\mu,0}\left[-\xi+\theta\eta|\mathcal{F}_{t}\right],\quad\forall t\in[0,T].$
Letting $\theta\rightarrow 1$, by [30, Theorem 3.2] and Lemma 2.3(vi), we have
$\rho_{t}(\xi)-\rho_{t}(\eta)\leqslant\mathcal{E}^{\mu,0}[\eta-\xi]\leqslant\mathcal{E}^{\mu,0}[|\xi-\eta|],\quad t\in[0,T].$
By interchanging $\xi$ and $\eta$, we have
$|\rho_{t}(\xi)-\rho_{t}(\eta)|\leqslant\mathcal{E}^{\mu,0}[|\xi-\eta|],\quad\forall t\in[0,T].$ (6.18)
For $\xi\in L^{2}(\mathcal{F}_{T})$, we set $\xi_{n}:=(\xi\vee(-n))\wedge n$. Then, by Step 1, we have $\rho_{t}(\xi_{n})=\mathcal{E}^{g}[-\xi_{n}|\mathcal{F}_{t}]$. This with (6.18) and [30, Theorem 3.2] implies that for all $\xi\in L^{2}(\mathcal{F}_{T})$, (3.2) holds.
From the same arguments as in Step 1, we get that for all $\xi\in L^{2}(\mathcal{F}_{T})$, (3.4) holds. The proof is complete. ∎
Appendix: The proof of (6.9)
Proof.
By the definition of $f$, we have, $dt\times dP$-$a.e.,$
$f(t,x)\geqslant\sup_{z\in\mathbf{R}^{d}}\{x\cdot z-\mu|z|-\nu|z|^{2}\}\geqslant\frac{(|x|-\mu)^{2}}{4\nu}1_{\{|x|>\mu\}}\geqslant\frac{1}{8\nu}(|x|^{2}-2\mu^{2}),\quad\forall x\in\mathbf{R}^{d}.$ (a.1)
By [12, Theorem 2.6(i)], for $\xi\in\mathcal{L}^{\exp}_{T},$ the following BSDE:
$Y_{t}=-\xi+\int_{t}^{T}g(s,Z_{s})ds-\int_{t}^{T}Z_{s}\cdot dB_{s},\quad t\in[0,T].$ (a.2)
has a unique solution $(Y_{t},Z_{t})\in\mathcal{L}^{\exp}_{\mathcal{F}}\times\bigcap_{r>1}\mathcal{H}_{d}^{r}.$ By measurable selection theorem, we can find a progressively measurable process $\tilde{q}_{t}\in\partial g(t,Z_{t})$, $dt\times dP$-$a.e.$ (see also [3, Lemma 7.5]). This implies that $dt\times dP\text{-}a.e.,$
$f(t,\tilde{q}_{t})=\tilde{q}_{t}\cdot Z_{t}-g(t,Z_{t}).$ (a.3)
By (a.1) and (a.3), we get that $dt\times dP\text{-}a.e.,$
$\frac{1}{8\nu}(|\tilde{q}_{t}|^{2}-2\mu^{2})\leqslant f(t,\tilde{q}_{t})$ $\leqslant|\tilde{q}_{t}\cdot Z_{t}|+|g(t,Z_{t})|$
$\leqslant 16\nu|Z_{t}|^{2}+\frac{|\tilde{q}_{t}|^{2}}{16\nu}+(\mu+\nu)(1+|Z_{t}|^{2}),$ (a.4)
where implies that $\tilde{q}_{t}\in H_{d}^{2}$.
For any $q_{t}\in H_{d}^{2}$, we set
$\theta_{t}^{q}:=\exp\left(\int_{0}^{t}q_{s}\cdot dB_{s}-\frac{1}{2}\int_{0}^{t}|q_{s}|^{2}ds\right),\quad\forall t\in[0,T].$
We define the following stopping time:
$\delta_{n}:=\inf\left\{s\geqslant 0:\int_{0}^{s}|\tilde{q}_{r}|^{2}dr\geqslant n\right\}\wedge T.$
and set $\tilde{q}_{t}^{n}:=\tilde{q}_{t}1_{\{t\in[0,\delta_{n}]\}}$. By [13, Proposition 4.3 and Inequality (4.6)], we have
$E[\theta^{\tilde{q}^{n}}_{T}\ln(1+\theta^{\tilde{q}^{n}}_{T})]\leqslant\frac{1}{2}E\left[\int_{0}^{T}\theta^{\tilde{q}^{n}}_{s}|\tilde{q}_{s}^{n}|^{2}ds\right]+\ln 2,$ (a.5)
and
$|x||y|\leqslant\beta\exp\left(\frac{|x|}{\beta}\right)+\beta|y|\ln(1+|y|),\quad\forall x,y\in\mathbf{R}^{d},\ \forall\beta>0.$ (a.6)
By (a.5), (a.1)-(a.3) and (a.6), we have
$\frac{1}{4\nu}(E[\theta^{\tilde{q}^{n}}_{T}\ln(1+\theta^{\tilde{q}^{n}}_{T})]-\ln 2)-\frac{\mu^{2}T}{4\nu}$ $\leqslant\frac{1}{8\nu}E\left[\int_{0}^{T}1_{\{s\in[0,\delta_{n}]\}}\theta^{\tilde{q}^{n}}_{s}(|\tilde{q}_{s}^{n}|^{2}-2\mu^{2})ds\right]$
$\leqslant E\left[\int_{0}^{T}1_{\{s\in[0,\delta_{n}]\}}\theta^{\tilde{q}^{n}}_{s}f(s,\tilde{q}_{s}^{n})ds\right]$
$\leqslant E\left[\int_{0}^{T}1_{\{s\in[0,\delta_{n}]\}}E[\theta^{\tilde{q}^{n}}_{T}|\mathcal{F}_{s}]({\tilde{q}_{s}^{n}}\cdot Z_{s}-g(s,Z_{s}))ds\right]$
$=E\left[\theta^{\tilde{q}^{n}}_{T}\int_{0}^{T}1_{\{s\in[0,\delta_{n}]\}}({\tilde{q}_{s}^{n}}\cdot Z_{s}-g(s,Z_{s}))ds\right]$
$=E\left[\theta^{\tilde{q}^{n}}_{T}\left(Y_{\delta_{n}}-Y_{0}-\int_{0}^{\delta_{n}}Z_{s}\cdot dB^{\tilde{q}^{n}}_{s}\right)\right]$
$\leqslant\frac{1}{8\nu}E\left[\sup_{t\in[0,T]}\exp(16\nu|Y_{t}|)\right]+\frac{1}{8\nu}E[\theta^{\tilde{q}^{n}}_{T}\ln(1+\theta^{\tilde{q}^{n}}_{T})],$ (a.7)
where $B^{\tilde{q}^{n}}_{t}:=B_{t}-\int_{0}^{t}{\tilde{q}^{n}}_{s}ds$ is a standard Brownian motion under the probability $Q^{\tilde{q}^{n}}$ satisfying $Q^{\tilde{q}^{n}}/dP=\theta^{\tilde{q}^{n}}_{T}$. It follows from (Appendix: The proof of (6.9)) that there exists a constant $C$ independent of $n$ such that $E[\theta^{\tilde{q}^{n}}_{T}\ln(1+\theta^{\tilde{q}^{n}}_{T})]\leqslant C.$ Then by Protter [31, Theorem 11], we get that $\{\theta^{\tilde{q}^{n}}_{T}\}_{n\geqslant 1}$ is uniformly integrable, and then by Vitali convergence theorem, we have $E[\theta^{\tilde{q}}_{T}]=1$, which implies that $\theta^{\tilde{q}}_{t}$ is a uniformly integrable martingale. Moreover, by Fatou’s lemma, we also have
$E[\theta^{\tilde{q}}_{T}\ln(1+\theta^{\tilde{q}}_{T})]\leqslant C.$ (a.8)
Set
$\tau_{n}^{t}:=\inf\left\{s\geqslant t:\int_{t}^{s}|Z_{r}|^{2}dr\geqslant n\right\}\wedge T.$
By (a.2) and (a.3), we have
$Y_{t}$ $=E_{Q^{\tilde{q}}}\left[Y_{\tau_{n}^{t}}+\int_{t}^{\tau_{n}^{t}}g(s,Z_{s})-\int_{t}^{\tau_{n}^{t}}Z_{s}\cdot dB_{s}|\mathcal{F}_{t}\right]$
$=E_{Q^{\tilde{q}}}\left[Y_{\tau_{n}^{t}}-\int_{t}^{\tau_{n}^{t}}f(s,\tilde{q}_{s})-\int_{t}^{\tau_{n}^{t}}Z_{s}\cdot dB^{\tilde{q}}_{s}|\mathcal{F}_{t}\right]$
$=E_{Q^{\tilde{q}}}\left[Y_{\tau_{n}^{t}}-\int_{t}^{\tau_{n}^{t}}f(s,\tilde{q}_{s})ds|\mathcal{F}_{t}\right],\quad\forall t\in[0,T],$ (a.9)
where $B^{\tilde{q}}_{t}:=B_{t}-\int_{0}^{t}{\tilde{q}}_{s}ds$ is standard Brownian motion under the probability $Q^{\tilde{q}}$. By (a.6) and (a.8), we have
[…]
By (Appendix: The proof of (6.9)), (a.10), Fatou’s lemma and the fact that $f\geqslant 0$, we have
$E_{Q^{\tilde{q}}}\left[\int_{t}^{T}|f(s,\tilde{q}_{s})|ds|\mathcal{F}_{t}\right]<\infty,\quad\forall t\in[0,T].$ (a.11)
From (a.10) and (a.11), it follows that $\tilde{q}_{t}\in\mathcal{Q}_{(\xi,f)}$. Then, by (Appendix: The proof of (6.9))-(a.11) and domination convergence theorem, we have
$Y_{t}=E_{Q^{\tilde{q}}}\left[-\xi-\int_{t}^{T}f(s,\tilde{q}_{s})ds|\mathcal{F}_{t}\right],\quad\forall t\in[0,T].$ (a.12)
We next show that for any $q_{t}\in\mathcal{Q}_{(\xi,f)},$
$Y_{t}\geqslant E_{Q^{q}}\left[-\xi-\int_{t}^{T}f(s,q_{s})ds|\mathcal{F}_{t}\right],\quad\forall t\in[0,T].$ (a.13)
By (a.5), (a.1), (a.3) and the definition of $\mathcal{Q}_{(\xi,f)},$ we have
$\frac{1}{4\nu}(E[\theta^{q}_{T}\ln(1+\theta^{q}_{T})]-\ln 2)-\frac{\mu^{2}T}{4\nu}$ $\leqslant\frac{1}{8\nu}E\left[\int_{0}^{T}\theta^{q}_{s}(|q_{s}|^{2}-2\mu^{2})ds\right]$
$=E\left[\int_{0}^{T}E[\theta^{q}_{T}|\mathcal{F}_{s}]f(s,q_{s})ds\right]$
$=E_{Q^{q}}\left[\int_{0}^{T}f(s,q_{s})ds\right]<\infty,$ (a.14)
Since $f(t,q_{t})\geqslant q_{t}\cdot Z_{t}-g(t,Z_{t}),$ by (Appendix: The proof of (6.9)) and similar arguments as in (Appendix: The proof of (6.9))-(a.10), we can obtain (a.13), which together with (a.12) and Definition 2.2(i) gives (6.9). ∎
Declarations
Competing interests The authors declare no competing interests

### v4

@@
 Compared to [10], our result holds for unbounded terminal variables, and the function $f$ can be determined by the Legendre-Fenchel transformation of the generator $g$ in (3.2).
-Furthermore, for coherent DRMs satisfying (1.3), we obtain a dual representation as in (1.5), where the terminal variable $\xi$ is unbounded and $\mathcal{Q}$ is characterized by the subdifferential $\partial g(t,0)$.
+Furthermore, for coherent DRMs satisfying (1.3), we obtain a dual representation as in (1.5), where the terminal variable $\xi$ is unbounded and $\mathcal{Q}_{1}$ is characterized by the subdifferential $\partial g(t,0)$.
 Our method relies on the representation (3.2) and therefore requires (1.3).
@@
 ${\cal{G}}^{\mu,\nu}:=\{g\in{\cal{G}}:$ $dt\times dP$-$a.e.,$ $\forall z\in\mathbf{R}^{\mathit{d}},\ |g(t,z)|\leqslant\mu|z|+\nu|z|^{2}\};$
-${\cal{G}}^{\mu,\nu}_{\theta}:=\{g\in{\cal{G}}:$ $dt\times dP$-$a.e.,$ $\forall z,\tilde{z}\in\mathbf{R}^{d}$, $\forall\theta\in(0,1)$, $g(t,z)-\theta g(t,\tilde{z})\leqslant\mu|z-\tilde{z}|+\frac{\nu}{1-\theta}|z-\tilde{z}|^{2}\}$;
+${\cal{G}}^{\mu,\nu}_{\theta}:=\{g\in{\cal{G}}:$ $dt\times dP$-$a.e.,$ $\forall z,\tilde{z}\in\mathbf{R}^{d}$, $\forall\theta\in(0,1)$, $g(t,z)-\theta g(t,\tilde{z})\leqslant\mu|z-\theta\tilde{z}|+\frac{\nu}{1-\theta}|z-\theta\tilde{z}|^{2}\}$;
 ${\cal{G}}_{conv}:=\{g\in{\cal{G}}:$ $dt\times dP$-$a.e.,$ $g(t,z)$ is convex in z};z\};
@@
 For $g\in{\cal{G}}^{\mu,\nu}\cap{\cal{G}}^{\mu,\nu}_{\theta}$, we have $dt\times dP$-$a.e.,$ for all $z,\tilde{z}\in\mathbf{R}^{d}$ and $\theta\in(0,1)$,
+$g(t,z~)\displaystyle$ $g(t,\tilde{z})$
 […]
-$\leqslant\mu|z-\tilde{z}|+\frac{\nu}{1-\theta}|z-\tilde{z}|^{2}+(1-\theta)(\mu+\nu)(1+|z|^{2}+|\tilde{z}|^{2}).$
-which implies that
-$\lim_{\tilde{z}\rightarrow z}|g(t,z)-g(t,\tilde{z})|\leqslant(1-\theta)(\mu+\nu)(1+2|z|^{2})\rightarrow 0,\quad\text{as}\ \theta\rightarrow 1.$
+$\leqslant\mu|z-\theta\tilde{z}|+\frac{\nu}{1-\theta}|z-\theta\tilde{z}|^{2}+(1-\theta)(\mu|\tilde{z}|+\nu|\tilde{z}|^{2})$
+$\leqslant\mu(|z-\theta\tilde{z}|+|\tilde{z}-\theta z|)+\frac{\nu}{1-\theta}(|z-\theta\tilde{z}|^{2}+|\tilde{z}-\theta z|^{2})+(1-\theta)(\mu(|\tilde{z}|+|z|)+\nu(|\tilde{z}|^{2}+|z|^{2})).$
+By interchanging $z$ and $\tilde{z}$, we have
+$\lim_{\tilde{z}\rightarrow z}|g(t,z)-g(t,\tilde{z})|\leqslant 4(1-\theta)(\mu|z|+\nu|z|^{2})\rightarrow 0,\quad\text{as}\ \theta\rightarrow 1.$
 It follows that for all $g\in{\cal{G}}^{\mu,\nu}\cap{\cal{G}}^{\mu,\nu}_{\theta}$, $dt\times dP$-$a.e.,$ $g(t,z)$ is continuous in $z$.
@@
 Proof.
-Let $Y_{t}$ be a $g^{\mu,\nu}$-submartingale. For any $0\leqslant s\leqslant t\leqslant T$, by Zheng et al. [39, Corollary 3.6], we have
-$Y_{s}\leqslant\mathcal{E}^{\mu,\nu}[Y_{t}|{\cal{F}}_{s}]=\frac{1}{2\nu}\ln(E_{Q}[\exp(2\nu Y_{t})|{\cal{F}}_{s}]).$
+Let $Y_{t}$ be a $g^{\mu,\nu}$-submartingale. By Zheng et al. [39, Corollary 3.6], we have
+$Y_{s}\leqslant\mathcal{E}^{\mu,\nu}[Y_{t}|{\cal{F}}_{s}]=\frac{1}{2\nu}\ln(E_{Q}[\exp(2\nu Y_{t})|{\cal{F}}_{s}]),\ \ \forall 0\leqslant s\leqslant t\leqslant T.$
 Here, $Q$ is a probability measure such that $dQ/dP=\exp\{\int_{0}^{T}b_{s}\cdot dB_{s}-\frac{1}{2}\int_{0}^{T}|b_{s}|^{2}ds\}$, where $b_{t}$ is a progressively measurable process such that $|b_{t}|\leqslant\mu.$ This implies that for any $0\leqslant s\leqslant t\leqslant T$, $\exp(2\nu Y_{s})\leqslant E_{Q}[\exp(2\nu Y_{t})|{\cal{F}}_{s}]$.
@@
 Then by the classical martingale theory, we get that $Y_{t}$ admits an RCLL modification.
-Let $Y_{t}$ be a $g^{-\mu,-\nu}$-supermartingale. Since for any $0\leqslant s\leqslant t\leqslant T$,
-$Y_{s}\leqslant-{\cal{E}}^{-\mu,-\nu}[Y_{t}|{\cal{F}}_{s}]={\cal{E}}^{\mu,\nu}[-Y_{t}|{\cal{F}}_{s}],$
-we get that $Y_{t}$ is a $g^{\mu,\nu}$-submartingale. Thus, $Y_{t}$ admits an RCLL modification. ∎
+If $Y_{t}$ is a $g^{-\mu,-\nu}$-supermartingale, then $Y_{s}\leqslant-{\cal{E}}^{-\mu,-\nu}[Y_{t}|{\cal{F}}_{s}]={\cal{E}}^{\mu,\nu}[-Y_{t}|{\cal{F}}_{s}]$ for any $0\leqslant s\leqslant t\leqslant T$.
+This implies that $Y_{t}$ is a $g^{\mu,\nu}$-submartingale.
+Thus $Y_{t}$ admits an RCLL modification.
+∎
 Remark 2.6.
@@
 $\exp(2\nu\mathcal{E}^{\mu,\nu}[Y_{t}|{\cal{F}}_{s}])=\mathcal{E}^{\mu,0}[\exp(2\nu Y_{t})|{\cal{F}}_{s}],\quad\forall s\in[0,t].$ (2.2)
-Since for any $0\leqslant s\leqslant t\leqslant T$, $\exp(2\nu Y_{s})\leqslant\exp(2\nu\mathcal{E}^{\mu,\nu}[Y_{t}|{\cal{F}}_{s}])$, by (2.2), we get that $\exp(2\nu Y_{t})$ is a $g^{\mu,0}$-submartingale.
+Since for all $s\in[0,t]$, $\exp(2\nu Y_{s})\leqslant\exp(2\nu\mathcal{E}^{\mu,\nu}[Y_{t}|{\cal{F}}_{s}])$, by (2.2), $\exp(2\nu Y_{t})$ is a $g^{\mu,0}$-submartingale.
 Then, by Peng [29, Theorem 3.3], there exists a pair $(z_{t},a_{t})\in H_{d}^{2}\times{\cal{A}}$, such that
@@
 Applying Itô’s formula to $\frac{1}{2\nu}\ln(y_{t})$, we have
-$Y_{t}=Y_{T}+$ $\int_{t+}^{T}\frac{1}{2\nu y_{s-}}\left(\mu|z_{s}|ds-da_{t}-z_{s}\cdot dB_{s}\right)+\int_{t+}^{T}\frac{|z_{s}|^{2}}{4\nu|y_{s}|^{2}}ds-\sum_{s\in(t,T]}\left(\Delta Y_{s}-\frac{1}{2\nu y_{s-}}\Delta y_{s}\right)$
+$Y_{t}=Y_{T}+$ $\int_{t+}^{T}\frac{1}{2\nu y_{s-}}\left(\mu|z_{s}|ds-da_{s}-z_{s}\cdot dB_{s}\right)+\int_{t+}^{T}\frac{|z_{s}|^{2}}{4\nu|y_{s}|^{2}}ds-\sum_{s\in(t,T]}\left(\Delta Y_{s}-\frac{1}{2\nu y_{s-}}\Delta y_{s}\right)$
 $=Y_{T}+$ $\int_{t}^{T}\left(\mu\left|\frac{z_{s}}{2\nu y_{s}}\right|+\nu\left|\frac{z_{s}}{2\nu y_{s}}\right|^{2}\right)ds-\int_{t}^{T}\frac{z_{s}}{2\nu y_{s}}\cdot dB_{s}$
@@
 is a convex DRM, called dynamic entropy risk measure. By Itô’s formula, we have
-$\rho^{\mathrm{ent}}(\xi)=\mathcal{E}^{0,\nu}[-\xi|\mathcal{F}_{t}],\quad t\in[0,T],$
+$\rho^{\mathrm{ent}}_{t}(\xi)=\mathcal{E}^{0,\nu}[-\xi|\mathcal{F}_{t}],\quad t\in[0,T],$
 and thus $\rho$ satisfies (A). From Theorem 3.3, it follows that
 $\rho^{\mathrm{ent}}_{t}(\xi)=\operatorname{ess\,sup}_{q\in\mathcal{Q}_{\xi,f}}E_{Q^{q}}\left[-\xi-\int_{t}^{T}\frac{|q_{s}|^{2}}{4\nu}ds|\mathcal{F}_{t}\right],\quad\forall t\in[0,T].$ (3.5)
-In particular, when $\xi\in L^{\infty}(\mathcal{F}_{T})$, by (Appendix: The proof of (6.9)) in the Appendix, $\mathcal{Q}_{\xi,f}$ in (3.4) can be reduced to $\mathcal{Q}:=\{Q^{q}:$ $dQ^{q}/dP=\theta^{q}_{T}:=\int_{0}^{T}q_{s}\cdot dB_{s}-\frac{1}{2}\int_{0}^{T}|q_{s}|^{2}ds$ with $q_{t}\in\mathcal{H}_{d}^{BMO}\}.$ For $Q^{q}\in\mathcal{Q}$, the conditional relative entropy of $Q^{q}$ with respect to $P$:
+In particular, when $\xi\in L^{\infty}(\mathcal{F}_{T})$, by (Proof.) in the Appendix, $\mathcal{Q}_{\xi,f}$ in (3.4) can be reduced to $\mathcal{Q}:=\{Q^{q}:$ $dQ^{q}/dP=\theta^{q}_{T}:=\int_{0}^{T}q_{s}\cdot dB_{s}-\frac{1}{2}\int_{0}^{T}|q_{s}|^{2}ds$ with $q_{t}\in\mathcal{H}_{d}^{BMO}\}.$ For $Q^{q}\in\mathcal{Q}$, the conditional relative entropy of $Q^{q}$ with respect to $P$:
 $\mathbb{H}_{t}(Q^{q}|P):=E_{Q^{q}}\left[\ln\left(\frac{\theta^{q}_{T}}{\theta^{q}_{t}}\right)|\mathcal{F}_{t}\right]$ $=E_{Q^{q}}\left[\left(\int_{t}^{T}q_{s}\cdot dB_{s}-\frac{1}{2}\int_{t}^{T}|q_{s}|^{2}ds\right)|\mathcal{F}_{t}\right]$
@@
 where $dt\times dP$-$a.e.,$ $|g(t,z)|\leqslant\mu|z|+\nu|z|^{2}$ and for all $\tilde{z}\in\textbf{R}^{d}$ and $\theta\in(0,1)$,
-$g(t,z)-\theta g(t,\tilde{z})\leqslant\mu|z-\tilde{z}|+\frac{\nu}{1-\theta}|z-\tilde{z}|^{2}.$
+$g(t,z)-\theta g(t,\tilde{z})\leqslant\mu|z-\theta\tilde{z}|+\frac{\nu}{1-\theta}|z-\theta\tilde{z}|^{2}.$
 This implies that $g\in{\cal{G}}^{\mu,\nu}\cap{\cal{G}}^{\mu,\nu}_{\theta}$.
@@
 Since $\rho$ satisfies (r5) and (r6) on $\mathcal{L}^{\exp}_{T}$, by Lemma 2.3(ix), we get that $g\in{\cal{G}}^{\mu,\nu}\cap{\cal{G}}_{subl}$.
-This implies that for all $z\in\mathbf{R}^{d}$ and $\beta>0$, $\beta|g(z)|=|g(\beta z)|\leqslant\beta\mu|z|+\beta^{2}\nu|z|^{2}$, and thus
-$|g(z)|\leqslant\mu|z|+\beta\nu|z|^{2}.$
+This implies that $dt\times dP$-$a.e.$, for all $z\in\mathbf{R}^{d}$ and $\beta>0$, $\beta|g(t,z)|=|g(t,\beta z)|\leqslant\beta\mu|z|+\beta^{2}\nu|z|^{2}$, and thus
+$|g(t,z)|\leqslant\mu|z|+\beta\nu|z|^{2}.$
 Letting $\beta\rightarrow 0$, we have $g\in{\cal{G}}^{\mu,0}\cap{\cal{G}}_{subl}$.
@@
 where $B^{\tilde{q}^{n}}_{t}:=B_{t}-\int_{0}^{t}{\tilde{q}^{n}}_{s}ds$ is a standard Brownian motion under the probability $Q^{\tilde{q}^{n}}$ satisfying $Q^{\tilde{q}^{n}}/dP=\theta^{\tilde{q}^{n}}_{T}$.
-It follows from (Appendix: The proof of (6.9)) that there exists a constant $C$ independent of $n$ such that $E[\theta^{\tilde{q}^{n}}_{T}\ln(1+\theta^{\tilde{q}^{n}}_{T})]\leqslant C.$ Then by Protter [31, Theorem 11], we get that $\{\theta^{\tilde{q}^{n}}_{T}\}_{n\geqslant 1}$ is uniformly integrable, and then by Vitali convergence theorem, we have $E[\theta^{\tilde{q}}_{T}]=1$, which implies that $\theta^{\tilde{q}}_{t}$ is a uniformly integrable martingale.
+It follows from (Proof.) that there exists a constant $C$ independent of $n$ such that $E[\theta^{\tilde{q}^{n}}_{T}\ln(1+\theta^{\tilde{q}^{n}}_{T})]\leqslant C.$ Then by Protter [31, Theorem 11], we get that $\{\theta^{\tilde{q}^{n}}_{T}\}_{n\geqslant 1}$ is uniformly integrable, and then by Vitali convergence theorem, we have $E[\theta^{\tilde{q}}_{T}]=1$, which implies that $\theta^{\tilde{q}}_{t}$ is a uniformly integrable martingale.
 Moreover, by Fatou’s lemma, we also have
@@
 […]
-By (Appendix: The proof of (6.9)), (a.10), Fatou’s lemma and the fact that $f\geqslant 0$, we have
+By (Proof.), (a.10), Fatou’s lemma and the fact that $f\geqslant 0$, we have
 $E_{Q^{\tilde{q}}}\left[\int_{t}^{T}|f(s,\tilde{q}_{s})|ds|\mathcal{F}_{t}\right]<\infty,\quad\forall t\in[0,T].$ (a.11)
-From (a.10) and (a.11), it follows that $\tilde{q}_{t}\in\mathcal{Q}_{(\xi,f)}$. Then, by (Appendix: The proof of (6.9))-(a.11) and domination convergence theorem, we have
+From (a.10) and (a.11), it follows that $\tilde{q}_{t}\in\mathcal{Q}_{(\xi,f)}$. Then, by (Proof.)-(a.11) and domination convergence theorem, we have
 $Y_{t}=E_{Q^{\tilde{q}}}\left[-\xi-\int_{t}^{T}f(s,\tilde{q}_{s})ds|\mathcal{F}_{t}\right],\quad\forall t\in[0,T].$ (a.12)
@@
 $=E_{Q^{q}}\left[\int_{0}^{T}f(s,q_{s})ds\right]<\infty,$ (a.14)
-Since $f(t,q_{t})\geqslant q_{t}\cdot Z_{t}-g(t,Z_{t}),$ by (Appendix: The proof of (6.9)) and similar arguments as in (Appendix: The proof of (6.9))-(a.10), we can obtain (a.13), which together with (a.12) and Definition 2.2(i) gives (6.9).
-∎
+Since $f(t,q_{t})\geqslant q_{t}\cdot Z_{t}-g(t,Z_{t}),$ by (Proof.) and similar arguments as in (Proof.)-(a.10), we can obtain (a.13), which together with (a.12) and Definition 2.2(i) gives (6.9). ∎
 Declarations